### v1

# CorrSteer: Steering Improves Task Performance and Safety in LLMs through Correlation-based Sparse Autoencoder Feature Selection

Abstract
Sparse Autoencoders (SAEs) can extract interpretable features from large language models (LLMs) without supervision. However, their effectiveness in downstream steering tasks is limited by the requirement for contrastive datasets or large activation storage. To address these limitations, we propose CorrSteer, which selects features by correlating sample correctness with SAE activations from generated tokens at inference time. This approach uses only inference-time activations to extract more relevant features, thereby avoiding spurious correlations. It also obtains steering coefficients from average activations, automating the entire pipeline. Our method shows improved task performance on QA, bias mitigation, jailbreaking prevention, and reasoning benchmarks on Gemma 2 2B and LLaMA 3.1 8B, notably achieving a +4.1% improvement in MMLU performance and a +22.9% improvement in HarmBench with only 4000 samples. Selected features demonstrate semantically meaningful patterns aligned with each task’s requirements, revealing the underlying capabilities that drive performance. Our work establishes correlation-based selection as an effective and scalable approach for automated SAE steering across language model applications.
CorrSteer: Steering Improves Task Performance and Safety in LLMs through Correlation-based Sparse Autoencoder Feature Selection
Seonglae Cho1,2  Zekun Wu1,2  Adriano Koshiyama1,2 1Holistic AI  2University College London
1 Introduction
Sparse Autoencoders (SAEs) have emerged as a powerful tool for decomposing superposed representations in large language models (LLMs) into interpretable sparse latent dimensions (Huben et al., 2023). By reconstructing neural activations through a sparse bottleneck, SAEs effectively disentangle semantic features that can be leveraged for downstream tasks such as probing and steering (Bricken et al., 2023).
However, existing SAE-based steering approaches face significant limitations: (1) contrastive datasets (Soo et al., 2025) or large activation storage (Zhao et al., 2025; Arad et al., 2025) are required to identify the direction of the steering, and (2) they rely on the hidden states of context tokens to select both the features and their coefficients.
Figure 1: Top correlated features with MMLU in each layer of Gemma 2 2B.
Consequently, current use cases of SAE-based steering have been restricted to specific applications, such as bias mitigation (Durmus et al., 2024), knowledge unlearning (Muhamed et al., 2025; Wang et al., 2025; Zhou et al., 2025; Cywiński and Deja, 2025), and jailbreaking prevention (O’Brien et al., 2025). Moreover, SAE feature selection in these applications does not directly reflect language models’ generation capabilities, potentially limiting their applicability.
Figure 2: Benchmark performance of CorrSteer variants compared with the baseline on Gemma 2 2B.
To address these limitations, this work introduces CorrSteer, which leverages generation-time features by correlating with task outcomes for task-specific feature selection and steering coefficient determination. Our approach employs Pearson correlation, which captures linear relationships, a lightweight yet effective criterion for rapidly identifying task-relevant features from minimal samples. Focusing on steering static behaviors, CorrSteer’s effectiveness is demonstrated on generation tasks by improving benchmark accuracy on MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024), BBQ (Parrish et al., 2022), HarmBench (Mazeika et al., 2024), XSTest (Röttger et al., 2024), and SimpleQA (Wei et al., 2024). Finally, by defining SER (Side Effect Ratio), three variants of CorrSteer are compared targeting the minimization of SER against fine-tuning.
2 Background
Mechanistic interpretability aims to reverse-engineer neural networks into human-interpretable components (Olah et al., 2020; Elhage et al., 2021). A central challenge in this endeavor is the superposition phenomenon, where neural networks learn to represent more features than available dimensions (Elhage et al., 2022). This efficient representation strategy complicates efforts to identify the consistent role of specific latent dimensions.
2.1 Sparse Autoencoders
Sparse Autoencoders (Huben et al., 2023; Bricken et al., 2023) address the superposition problem by learning to decompose neural activations into interpretable, sparse features. Given an activation vector $\mathbf{x}\in\mathbb{R}^{d}$, an SAE learns an encoder $f_{\text{enc}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ and decoder $f_{\text{dec}}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ where $k\gg d$, such that:
$\mathbf{z}$ $=f_{\text{enc}}(\mathbf{x})=\text{Activation}(\mathbf{W}_{\text{enc}}\mathbf{x}+\mathbf{b}_{\text{enc}})$ (1)
$\hat{\mathbf{x}}$ $=f_{\text{dec}}(\mathbf{z})=\mathbf{W}_{\text{dec}}\mathbf{z}+\mathbf{b}_{\text{dec}}$ (2)
The training objective is usually a combination of reconstruction loss with sparsity regularization:
$\mathcal{L}=\|\mathbf{x}-\hat{\mathbf{x}}\|^{2}+\lambda\|\mathbf{z}\|_{1}$ (3)
2.2 Steering Vectors
Steering vectors (Subramani et al., 2022) represent a class of methods for controlling neural network outputs by manipulating internal activations. Traditional approaches, such as CAA (Rimsky et al., 2024), compute activation differences between contrasting examples and apply these differences. For precise control to not inadvertently affect related behaviors simultaneously, PaCE (Luo et al., 2024) utilizes sparse coding for orthogonal steering directions.
2.3 SAE-based Steering
SAE-based steering leverages SAE latents for predictable control based on feature semantics. SAE-TS (Chalnev et al., 2024; Soo et al., 2025) reduces the side effects of steering by linearly approximating feature directions. SPARE (Zhao et al., 2025) utilizes Mutual Information to select features and their coefficients but requires large activation storage due to its non-linearity. DSG (Muhamed et al., 2025) utilizes Fisher Information Matrix to select features but requires contrastive datasets and additional backward computation. Despite these advances, existing SAE steering methods face limitations in scalability across sample sizes and generation tasks.
3 Method
Linear correlation offers both interpretability and faithfulness as a criterion for feature selection. SAEs capture linear relationships, consistent with the Linear Representation Hypothesis (Socher et al., 2013; Faruqui et al., 2015; Park et al., 2023), and have a proven ability to disentangle interpretable features in a linear manner. The faithfulness of Pearson correlation is further supported by recent work from Oikarinen et al. (2025).
3.1 Correlation-based Feature Selection
Our approach, CorrSteer, centers on the observation that features most correlated with task performance are likely to be relevant for steering. The approach employs the Pearson correlation coefficient, applied only to generation-time features—specifically to the last token at each step.
Given a set of SAE features $\mathbf{z}=[z_{1},z_{2},\ldots,z_{D}]$ and corresponding task performance scores $\mathbf{y}=[y_{1},y_{2},\ldots,y_{n}]$ for $n$ samples, the correlation for each feature $i$ is computed as:
$r_{i}=\frac{\text{Cov}(z_{i},y)}{\sqrt{\text{Var}(z_{i})\cdot\text{Var}(y)}}$ (4)
To handle the computational challenges of large SAE feature dictionaries (typically $10^{4}$$10^{5}$ features), a streaming correlation accumulator is implemented that maintains $O(1)$ memory complexity:
Algorithm 1 Streaming Correlation Computation
 Initialize: $\sum x_{i}=0,\sum x_{i}^{2}=0,\sum x_{i}y_{i}=0,\sum y_{i}=0,\sum y_{i}^{2}=0,n=0$
for each batch $(\mathbf{X}_{\text{batch}},\mathbf{y}_{\text{batch}})$ do
  Update running sums for each feature dimension
  $n\leftarrow n+\text{batch\_size}$
end for
 Compute correlations:
$medmath{r_{i}=\frac{n\sum x_{i}y_{i}-\sum x_{i}\sum y_{i}}{\sqrt{(n\sum x_{i}^{2}-(\sum x_{i})^{2})(n\sum y_{i}^{2}-(\sum y_{i})^{2})}}}$
This computation is $O(1)$ with respect to sample size, and $O(LD)$ for fixed layer count $L$ and SAE latent dimension $D$. For generation tasks requiring multiple tokens, max-pooling is employed over valid token positions to aggregate feature activations, as empirically validated in our pooling comparison study (Table 5).
3.2 Coefficient Calculation
The steering coefficient for a selected feature $i$ is computed as the mean activation value among samples where task performance is positive. This approach is preferable to contrastive-based calculation, since SAE features produce non-negative activations due to their ReLU-based activation functions (Bricken et al., 2023) and thus cannot be meaningfully subtracted in a contrastive manner; negative activations are often unrelated noise (Joseph Bloom, 2024).
$c_{i}=\frac{1}{|\{j:y_{j}>0\}|}\sum_{j:y_{j}>0}z_{i,j}$ (5)
This ensures that the steering magnitude reflects the natural activation scale of the feature during successful task performance.
3.3 Steering Implementation
During inference, steering is applied by modifying residual stream activations. For a selected feature $i$, with coefficient $c_{i}$ and SAE decoder weights $\mathbf{W}_{\text{dec}}$ (the feature direction (Templeton et al., 2024)), the steering vector is:
$\mathbf{v}_{\text{steer}}=c_{i}\cdot\mathbf{W}_{\text{dec}}[:,i]$ (6)
The modified activation is:
$\mathbf{x}^{\prime}=\mathbf{x}+\mathbf{v}_{\text{steer}}$ (7)
Steering is applied to tokens corresponding to the positions from which the features were originally extracted, rather than only to the last token (Luo et al., 2024; Rimsky et al., 2024) or every token (Soo et al., 2025).
3.4 Feature Extraction Strategies
Figure 3: Comparison of selected features between CorrSteer-1, CorrSteer-A, and CorrSteer-P on the BBQ disambiguous task across all layers of Gemma 2 2B. Red points indicate features selected by each method.
For each layer $\ell$, we obtain SAE activations from the residual stream and rank features by correlation with task performance. The method compares a global view aggregated across layers and a layer-wise view for selecting features to steer. Three strategies are implemented:
  • CorrSteer-1: Select the single highest-correlated feature from the global view, allowing cross-layer feature competition.
  • CorrSteer-A: Select the top-1 feature within each layer.
  • CorrSteer-P: Prune features from CorrSteer-A using a validation set, retaining only those that improve baseline performance to reduce side effects from spurious correlations.
Each method has distinct pros and cons, discussed in Section 5. Only positively correlated features are selected to ensure steering induces positive activation, as our ablation study confirms that negative correlation features consistently degrade performance 6. All methods are fully automated based on observed activations without hyperparameter tuning.
3.5 Side Effect Ratio
A key challenge in correlation-based feature selection is distinguishing features that causally contribute to task success from those that merely correlate due to the model’s internal state, potentially causing unintended side effects. To quantify side effects, the Side Effect Ratio (SER) is defined as the proportion of negatively changed answers among all changed answers:
$SER=\frac{\text{\# negatively changed answers}}{\text{\# all changed answers}}$ (8)
This measure does not isolate the side effect of each individual feature; rather, it serves as a combined metric reflecting how well the selected features are optimized for the task without degrading the model’s original abilities. To reduce side effects, the approach focuses on features activated during the generation process, under the hypothesis that generation-time activations are more likely to be causally relevant to output. This inference-time focus is empirically validated by our pooling experiments (Table 5). Additionally, in the multi-layer approach, a validation-based filtering mechanism is introduced (CorrSteer-P), retaining only features that demonstrate actual steering effectiveness.
4 Experiments
Figure 4: Relation between consumed sample counts and test performance, final matched count of the selected features, and the most correlated features from each layer of Gemma 2 2B. The dotted lines show the baseline of default LLM performance and constrained decoding performance on MMLU answer options.
4.1 Experimental Setup
CorrSteer is evaluated across diverse generation benchmarks to demonstrate practical effectiveness.
Models and SAEs: Experiments are conducted using Gemma-2 2B (Team, 2024a) and LLaMA-3.1 8B (Team, 2024b) models, paired with their corresponding SAE releases from Gemma Scope (Lieberum et al., 2024) and LLaMA Scope (He et al., 2024), respectively. Both SAE families employ JumpReLU activation (Rajamanoharan et al., 2024). Additionally, the Gemma-2-2B-IT model with SAEs is employed, leveraging the fact that SAEs are typically transferable across fine-tuned models (Kissane et al., 2024), with proven low loss reported in the Gemma Scope paper (Lieberum et al., 2024).
Datasets: Our evaluation encompasses multiple benchmark categories:
  • Question Answering: MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024)
  • Reasoning: GSM8k (Cobbe et al., 2021)
  • Bias Evaluation: BBQ (Parrish et al., 2022)
  • Factuality: SimpleQA (Wei et al., 2024)
  • Safety: HarmBench (Mazeika et al., 2024), XSTest (Röttger et al., 2024)
For safety benchmarks, both the refusal benchmark HarmBench and the overrefusal benchmark XSTest are included to evaluate not only the steering ability for rejection but also the model’s capability to identify the context of requests.
Evaluation Metrics: For multiple-choice tasks (MMLU, MMLU-Pro, BBQ), exact match accuracy is used. For safety benchmarks, 1 - ASR (Attack Success Rate) is computed using a small refusal-detection language model. SimpleQA performance is measured using a small STS language model to match the expected answer, with more details in Appendix A.2.
A standard train-validation-test split is used for the CorrSteer pipeline. The training dataset is used to extract correlated SAE features, and the validation dataset is used to filter the most correlated features. The test dataset is used to evaluate the performance of the CorrSteer pipeline. Detailed configurations are provided in Appendix A.1.
Pooling Strategy: Two pooling strategies are available for coefficient and correlation calculations: max-pooling and mean-pooling. For multi-token generation tasks, max-pooling consistently outperforms mean-pooling, as empirically demonstrated in Table 5, likely due to its better capture of peak feature activations relevant to task success. However, for coefficient calculation in longer generation tasks such as GSM8K reasoning, mean-pooling is preferred as max-pooling produces excessively large coefficient values. Applying these large coefficients to every generated token degrades performance, leading to the adoption of mean-pooling for reasoning tasks.
4.2 SER Comparison
This study aims to demonstrate that CorrSteer can improve benchmark performance while maintaining low side effects. Using the defined SER, CorrSteer’s SER is compared against fine-tuning for a question-answering dataset. Additionally, the SER is compared when allowing or enforcing negatively correlated features, supporting the claim of SAE’s positive-tinted behavior. Finally, the SER is compared when pooling on every token rather than the inference-time generated token.
4.3 Feature Interpretability Analysis
The analysis is enhanced by incorporating feature descriptions from Neuronpedia, providing semantic interpretations of selected features. Safe/unsafe tendency analysis is conducted, along with task-wise and feature-level interpretations. This interpretability analysis validates that our correlation-based selection identifies semantically meaningful and task-relevant features, supporting the causal hypothesis underlying our approach.
Table 1: Performance comparison between baseline and CorrSteer variants across BBQ, MMLU, MMLU-Pro, GSM8K, SimpleQA, and XSTest on Gemma 2 2B. Results show accuracy (%) for all tasks.
Task Baseline CorrSteer-1 CorrSteer-P CorrSteer-A Fine-tuning
BBQ Ambig 59.10 62.38 66.65 62.08 -
BBQ Disambig 75.42 75.70 77.04 76.53 -
HarmBench 44.64 46.07 51.79 67.50 -
MMLU 52.23 52.82 55.83 56.32 55.85
MMLU-Pro 30.30 30.44 30.82 31.01 33.16
GSM8K 54.74 54.44 53.53 54.44 47.38
SimpleQA 3.63 3.76 3.96 3.80 -
XSTest 86.35 86.67 86.03 87.30 -
Table 2: Performance comparison between baseline and CorrSteer variants across BBQ, MMLU, MMLU-Pro, HarmBench, SimpleQA, and XSTest on LLaMA 3.1 8B. Results show accuracy (%) for all tasks.
Task Baseline CorrSteer-1 CorrSteer-P CorrSteer-A
BBQ Ambig 83.97 83.98 87.10 86.83
BBQ Disambig 90.07 90.13 90.33 90.30
HarmBench 0.71 0.36 15.71 17.86
MMLU 61.41 61.51 61.73 61.71
MMLU-Pro 32.13 32.55 35.08 34.71
SimpleQA 0.43 0.51 0.43 0.43
XSTest 61.27 62.22 62.22 58.41
5 Results
5.1 Generation Benchmark Performance
Table 1 and Table 2 present comprehensive results across our evaluation benchmarks. CorrSteer demonstrates consistent improvements across most tasks, including question answering, bias mitigation, and safety benchmarks, with the exception of the GSM8K reasoning task. This pattern suggests that our method enhances model adherence to static task requirements without introducing dynamic behaviors.
Method Comparison
In most cases, CorrSteer-A and CorrSteer-P demonstrate the highest performance, with CorrSteer-P showing particular dominance in LLaMA 3.1 8B. This observation is attributed to the less disentangled nature of Llama Scope features from superposition, which necessitates more aggressive pruning.
Safety Benchmarks:
For HarmBench, selected features enhance the model’s ability to refuse harmful requests, achieving a 22.9% improvement. However, for XSTest, improvement is limited to 1% due to the overrefusal characteristics of the benchmark. This is an expected result due to the static nature of CorrSteer, which hinders the ability to clearly distinguish between benign and harmful requests.
Table 3: Safety Evaluation Rate (SER) analysis for CorrSteer variants across different benchmarks on Gemma 2 2B. SER values closer to 0 indicate better safety performance.
CorrSteer-1 CorrSteer-P CorrSteer-A Fine-tuning
Task SER neg pos SER neg pos SER neg pos SER neg pos
BBQ Ambig 0.000 0 658 0.000 0 1532 0.076 53 649 - - -
BBQ Disambig 0.167 45 59 0.153 111 164 0.257 65 112 - - -
HarmBench 0.250 2 6 0.143 4 24 0.043 3 67 - - -
MMLU 0.355 11 20 0.172 249 286 0.202 264 299 0.407 1108 1616
MMLU-Pro 0.421 8 11 0.423 30 41 0.419 39 54 0.461 357 418
GSM8K 0.556 20 16 0.674 31 15 0.516 63 59 0.647 213 116
SimpleQA 0.167 1 5 0.188 3 13 0.353 6 11 - - -
XSTest 0.333 7 10 0.520 7 10 0.467 14 5 - - -
Table 4: Safety Evaluation Rate (SER) analysis for CorrSteer variants on Llama models across different benchmarks. SER values closer to 0 indicate better safety performance.
CorrSteer-1 CorrSteer-P CorrSteer-A
Task SER neg pos SER neg pos SER neg pos
BBQ Ambig 0.496 141 143 0.017 11 651 0.025 15 599
BBQ Disambig 0.433 45 59 0.404 111 164 0.367 65 112
HarmBench 0.333 3 6 0.226 7 24 0.171 6 29
MMLU 0.488 118 124 0.465 249 286 0.469 264 299
MMLU-Pro 0.355 11 20 0.280 40 103 0.310 45 100
SimpleQA 0.000 0 1 - 0 0 0.500 4 4
XSTest 0.412 7 10 0.412 7 10 0.737 14 5
Observations:
Results for both Gemma-2 2B and LLaMA 3.1 8B, presented in Table 1 and Table 2, demonstrate consistent patterns, with CorrSteer variants showing systematic improvements.
Multi-layer Superiority: Simultaneous multi-layer steering approaches, such as CorrSteer-A and CorrSteer-P, consistently outperform the single-layer CorrSteer-1 approach across all benchmarks, consistent with findings from other researchers (Liu et al., 2024; Zhao et al., 2025). This observation suggests that enabling features from different layers to compete globally yields superior task-relevant selections compared to layer-wise optimization.
Limited Factuality Impact: SimpleQA shows minimal improvement, confirming that CorrSteer enhances task adherence without introducing external factual knowledge. This is a desirable property, as it indicates the method improves model behavior rather than injecting information not present in the original model.
Comparison with Fine-tuning: CorrSteer demonstrates competitive performance while maintaining significantly lower side effect rates compared to fine-tuning. On MMLU, CorrSteer-A achieves higher accuracy (56.32%) than fine-tuning (55.85%) with substantially lower SER (0.202 vs 0.407). Similarly, on GSM8K, CorrSteer variants outperform fine-tuning in accuracy while maintaining lower side effect rates across all tasks.
5.2 Feature Analysis
Selected features demonstrate varying degrees of alignment with task requirements: while some benchmarks show consistent alignment through structured output features for multiple-choice tasks, refusal-related features for safety benchmarks, and task-specific semantic features for domain-specific evaluations, others exhibit less consistent patterns. The interpretability of selected features, validated through Neuronpedia descriptions, provides confidence in the semantic relevance of our correlation-based selection process. Notably, feature activation frequencies vary significantly across tasks, with performance improvements correlating with distinct activation patterns (Appendix 8). Analysis of selected features reveals semantically meaningful patterns aligned with task requirements:
Safety-related Features: HarmBench and BBQ demonstrate substantial improvements through selected safety-related features that enhance neutrality and refusal behavior. This interpretability analysis is covered in detail in Section 6.1.
5.3 SER Analysis
Table 3 presents the Safety Evaluation Rate analysis, demonstrating CorrSteer’s ability to minimize side effects while improving task performance.
Figure 5: SER comparison between different CorrSteer variants for Gemma 2 2B.
6 Discussion
This work establishes a viable and efficient approach for SAE-based steering, providing effective control across diverse applications. The interpretable feature combinations that yield performance improvements support the hypothesis that linear correlation serves as a meaningful unit for interpretable AI capabilities.
6.1 Feature Inspection
g A notable finding is that features selected by CorrSteer demonstrate task-relevant patterns that align with theoretical expectations Math-related features are consistently discovered across all tasks, proving beneficial even for bias mitigation and safety tasks. This universal correlation between mathematical features and accuracy aligns with DeepSeekMath (Shao et al., 2024)’s findings, where further pre-training on math-focused corpora yielded performance improvements across diverse tasks.
For BBQ features in LLaMA 3.1 8B, we observe:
  • L15/25166 themes of neutrality and balance in discourse (coeff: 0.259, corr: 0.433)
  • L25/10753 expressions of perception or belief in social dynamics (coeff: 1.147, corr: 0.428)
While bias-related features were expected for the BBQ benchmark, these neutrality-focused features demonstrate high positive correlation. Conversely, explicit bias-related and choice-making features exhibit negative correlations:
  • L8/8123 questions that ask for truthfulness or correctness regarding options or statements (coeff: 3.725, corr: -0.133)
  • L17/9134 choice-related phrases and expressions of preference (coeff: 2.379, corr: -0.451)
  • L19/15745 phrases related to decision-making and choice, particularly in the context of parenting and social interactions (coeff: 9.740, corr: -0.464)
These findings suggest that task-specific induced features contribute more to sample accuracy than meta-cognitive recognition features. Our ablation study further demonstrates that SAE-based sparse feature selection consistently outperforms raw activation steering across all evaluated tasks (Table 7).
6.2 Feature Set Transferability
The transferability of CorrSteer feature sets is evaluated across MMLU, MMLU-Pro, and BBQ benchmarks. Interestingly, our cross-task experiments reveal that MMLU features outperform task-specific features on BBQ Ambig and achieve comparable performance on MMLU-Pro, suggesting that some feature sets capture more generalizable reasoning patterns (Table 8). This transferability is attributed to their shared multiple-choice format, which requires similar structural feature patterns.
6.3 Task-Level Circuit
CorrSteer’s multi-layer approach relates to neural network circuit discovery research (Olah et al., 2020; Elhage et al., 2021). While emerging works focus on discovering task-specific circuits (Conmy et al., 2023; Marks et al., 2025; Ameisen et al., 2025; Lindsey et al., 2025; Sun, 2025), our steering vectors that work simultaneously across layers can be conceptualized as additive subgraphs of optimized task circuits, though they lack explicit interpretation of interactions and causality.
6.4 Side Effect Ratio
The primary challenge in AI steering concerns robustness for industrial applications, necessitating precise control mechanisms. Direct steering at each layer without updating original parameters based on token prediction distributions represents a key approach to minimize side effects. Theoretically, separating steering vectors across different activation spaces minimizes mutual interference in superpositioned states (Elhage et al., 2022).
6.5 Pooling Strategy Analysis
The results reveal interesting patterns across different pooling strategies (Table 5). Mean-pooling shows severe degradation on multi-token generation tasks (HarmBench: 0.00%, XSTest: 53.65%) where responses require multiple tokens. All-token pooling shows degraded performance on every task compared to max-pooling’s inference-time aggregation. This suggests that max-pooling better captures the critical activations needed for effective steering across all task types, while all-token pooling may introduce noise by including irrelevant token positions, and mean-pooling dilutes important signals by averaging across all tokens in longer sequences.
6.6 Computational Efficiency
Our streaming correlation implementation achieves O(1) memory complexity with respect to training set size, making the approach scalable to large datasets. The pipeline exhibits computational efficiency with minimal sample requirements (200-400) as demonstrated in Figure 4 and completes feature extraction within minutes. Static feature sets and coefficients at inference time eliminate SAE dependency during deployment.
6.7 Practical Applicability
CorrSteer operates as an auxiliary mechanism that captures correlations difficult to detect through supervised fine-tuning, applicable with minimal side effects even after fine-tuning. The automated pipeline enables rapid deployment across tasks and domains without hyperparameter tuning. The method shows effectiveness across two model families, as demonstrated on both Gemma-2 2B and LLaMA-3.1 8B models. Importantly, our approach has broader implications for AI safety, as demonstrated by its effectiveness in both bias mitigation and amplification (Table 9), highlighting the need for responsible deployment of such steering capabilities.
7 Conclusion
This work introduces CorrSteer, a fully automated pipeline for SAE-based language model steering that leverages correlation analysis. Our approach addresses key limitations of existing SAE steering methods by identifying task-relevant features without requiring manual feature exploration or contrastive datasets. Experimental validation across diverse benchmarks demonstrates CorrSteer’s effectiveness, consistently improving performance on question answering, bias mitigation, and safety evaluation tasks. Selected features for safety, mathematical, and refusal-related tasks reveal the underlying objectives and required capabilities that drive task performance.
Future Work
Several promising directions emerge from this work: Prompt Engineering Comparison: Future studies should compare CorrSteer with prompt engineering approaches, as prompt-based methods are expected to exhibit higher SER due to their less targeted intervention mechanisms. Dynamic Steering for Reasoning: The performance degradation observed in GSM8K reasoning tasks suggests the need for dynamic steering approaches that can adapt to the sequential nature of mathematical problem-solving, moving beyond static feature interventions. Orthogonal Feature Projection: To further minimize side effects, future work could explore feature filtering techniques that project out components already activated in baseline features before applying steering, potentially reducing interference with existing model capabilities.
Acknowledgments
The authors thank the teams behind Gemma Scope and LLaMA Scope for providing high-quality SAE releases that enabled this research. The authors also acknowledge Neuronpedia for providing automated feature descriptions that enhanced our interpretability analysis.
Limitations
The fundamental limitation of steering vectors is their static nature, which prevents adaptation to dynamic model behaviors. This constraint particularly affects reasoning tasks like GSM8K, where static steering cannot adequately handle the sequential nature of mathematical problem-solving. Our evaluation focuses primarily on discriminative and short-form generation tasks; long-form generation and creative tasks may require different approaches or modifications to our method.
Appendix A Appendix
A.1 Implementation Details
Feature Extraction:
Feature selection employs 4,000 samples across all datasets. For fair comparison, the same samples are used for training fine-tuning baselines. When datasets contain fewer than 4,000 samples, we use all available data. For datasets without predefined train/validation/test splits, we allocate 27% for training, 3% for validation, and 70% for testing. GSM8K uses 1,000 samples for feature selection with 50 samples reserved for validation.
Feature Steering:
Steering interventions are applied at the pre-execution stage of each transformer layer. The first layer is excluded from steering as the token embedding layer predominantly contains spurious correlations unrelated to the target tasks.
Fine-tuning
Fine-tuning is performed using AdamW optimizer with learning rate 1e-5 (reduced to 5e-6 for small datasets <2000 samples), weight decay 0.01, and gradient clipping at norm 1.0. The training schedule includes 3% warmup steps followed by cosine annealing decay. Training proceeds for one epoch with 4,000 samples, using exact target supervision where prompt tokens are masked with -100 labels and only target spans contribute to the loss.
A.2 Generation Benchmark Results
Evaluation Models: Two specialized models are employed for evaluation. The DistillRoBERTa model [^1] https://huggingface.co/datasets/huggingface-community/distill-roberta-toxicity-r1 is used to identify the rejection of harmful requests, while the BERT STS model [^2] https://huggingface.co/datasets/HuggingFaceH4/stsb_multi_mt is used for matching generated answers against expected responses.
A.3 Additional Results
Figure 6: Benchmark performance of CorrSteer variants compared with the default baseline on LLaMA 3.1 8B.
Task-Specific Analysis
MMLU: The global method selects features related to structured output formatting, addressing Gemma-2’s tendency to generate tokens outside the required A/B/C/D options. Post-steering, this hallucination issue is largely resolved.
MMLU-Pro: A similar issue occurs more severely due to the 10 options in MMLU-Pro. Constrained decoding, which samples tokens exclusively from available options, is applied to improve the model’s authentic capability, resulting in performance that remains higher than baseline, with CorrSteer-A achieving maximum performance.
BBQ: Similar improvements in format adherence are observed, with selected features promoting appropriate response structure.
Figure 7: SER comparison across datasets between different CorrSteer variants on LLaMA 3.1 8B.
Feature Frequency Analysis
We observe a strong correlation between feature activation frequency and CorrSteer’s performance improvements across tasks. As demonstrated in Figure 8, HarmBench exhibits consistently high activation frequencies across all layers, while SimpleQA shows frequencies approaching zero.
This pattern contrasts with the typical sparse activation nature of SAE features, where low frequency activation (below 5%) is considered normal and interpretable, while higher frequencies typically indicate non-interpretable (Stolfo et al., 2025; Smith et al., 2025). However, discovering task-specific features with near-100% activation frequency suggests these features are deeply related to the task requirements, resulting in substantial performance improvements for such tasks. Even for tasks with lower feature frequencies, CorrSteer maintains its advantage by preserving low SER values.
Figure 8: Frequency of activation samples across layers of Gemma 2 2B for SimpleQA (left) and HarmBench (right) tasks.
A.4 Ablation Study
Pooling Strategy
For generation tasks requiring multiple tokens, max-pooling is employed over valid token positions to aggregate feature activations before correlation computation. Our comprehensive evaluation confirms max-pooling’s superiority over alternative strategies (Table 5). However, for coefficient calculation in longer generation tasks such as GSM8K reasoning, mean-pooling is preferred as max-pooling produces excessively large coefficient values that degrade performance when applied to every generated token.
We also evaluate alternative pooling strategies including mean-pooling and all-token pooling for feature activation aggregation. The comparison results are presented in Table 5.
Table 5: Pooling strategy comparison on Gemma 2 2B using CorrSteer-A. For single-token generation such as MMLU, MMLU-Pro, and BBQ, mean-pooling naturally achieves identical performance to max-pooling since only one token is generated, while all-token pooling shows degraded performance. Mean-pooling shows severe degradation on multi-token generation tasks, demonstrating the superiority of max-pooling.
Task Baseline Max-pooling Mean-pooling All-token
MMLU 52.23 56.32 56.32 52.91
MMLU-Pro 30.30 31.00 31.00 30.16
BBQ Disambig 75.42 76.53 76.53 75.00
BBQ Ambig 59.10 62.08 62.08 57.98
HarmBench 44.64 67.50 0.00 47.14
XSTest 86.35 87.30 53.65 86.35
SimpleQA 3.63 3.80 3.76 3.73
Negative Correlation Features
To validate our design choice of using only positively correlated features, we conduct ablation experiments using negatively correlated features for steering. We compare two approaches: single-layer negative steering (CorrSteer-1 with negative features) and multi-layer negative steering (CorrSteer-A with negative features).
Table 6: Performance comparison between positive and negative correlation feature steering on Gemma 2 2B. Negative correlation features consistently show poor performance, validating our positive-only approach.
Task Baseline CorrSteer-A Negative-1 Negative-G
MMLU 52.23 56.32 52.24 49.45
MMLU-Pro 14.00 17.56 14.24 0.66
BBQ Disambig 75.42 76.53 75.37 12.15
BBQ Ambig 59.10 62.08 59.22 60.85
HarmBench 44.64 67.50 44.64 47.86
XSTest 86.35 87.30 86.35 86.67
SimpleQA 3.63 3.80 3.76 3.76
The results demonstrate that negative correlation features provide minimal improvement in single-layer steering and often cause severe performance degradation in multi-layer steering. Notably, MMLU-Pro drops to 0.66% and BBQ Disambig to 12.15% with negative multi-layer steering, confirming that negative correlations often represent spurious patterns rather than causal relationships. Additionally, combining positive and negative features simultaneously yields inferior performance compared to positive-only selection. This validates our approach of using only positively correlated features, which aligns with the non-negative nature of SAE activations.
Table 7: Performance comparison between raw activation steering and SAE-decoded steering on Gemma 2 2B. Decoding adds SAE decoder bias term for the first layer, while Decoding-A adds multi-layer feature directions as CorrSteer-A.
Task Baseline Raw Activation Decoding-1 Decoding-A CorrSteer-A
MMLU 52.23 49.85 55.38 54.38 56.32
MMLU-Pro 30.30 27.17 29.79 29.93 31.00
BBQ Disambig 75.42 75.71 77.00 75.03 76.53
BBQ Ambig 59.10 58.42 54.00 55.76 62.08
Raw Activation Steering
To validate the effectiveness of SAE-based sparse feature selection, we compare steering performance using raw residual stream activations. The results demonstrate a clear performance hierarchy: CorrSteer-A > SAE Decoding > Raw Activation across all evaluated tasks, which is explainable by Superposition Hypothesis (Elhage et al., 2022). One exception occurred in BBQ Disambig, where Decoding-1 shows better performance than CorrSteer-A. However, Decoding-1 failed to show robustness across benchmarks, frequently degrading performance while CorrSteer-A shows consistent performance across all tasks.
SAE Decoder Bias
Adding SAE decoder bias terms alongside selected features improves performance only at single-token generation tasks (BBQ, MMLU, MMLU-Pro). This effect appears related to attention sink mechanisms (Xiao et al., 2024), where increased residual stream norms amplify attention patterns in subsequent layers, acting similar to "response prefix" (Hazra et al., 2025). For constrained generation tasks, this norm amplification reduces hallucination by strengthening adherence to output format constraints. However, this enhancement is incompatible with multi-layer steering and diminishes when applied across multiple layers or tokens, with excessive application potentially causing model collapse.
A.5 Cross-Task Feature Transferability
To evaluate the transferability of selected features across different tasks, we conduct cross-task steering experiments where features selected for one task are applied to different target tasks. This analysis provides insights into the generalizability of task-specific feature sets.
Table 8: Cross-task feature transferability results on Gemma 2 2B. Features selected from source tasks (rows) are applied to target tasks (columns). Results show accuracy (%) with baseline performance in parentheses. MMLU-Pro results do not use constrained decoding, achieving 17.56% compared to unconstrained baseline (14.00%).
Source → Target MMLU MMLU-Pro BBQ Disambig BBQ Ambig
MMLU 56.32 (52.23) 19.67 (14.00) 74.62 (75.42) 64.01 (59.10)
MMLU-Pro 55.73 (52.23) 17.56 (14.00) 76.10 (75.42) 60.97 (59.10)
BBQ Disambig 54.74 (52.23) 16.11 (14.00) 76.53 (75.42) 60.85 (59.10)
BBQ Ambig 53.85 (52.23) 11.01 (14.00) 76.10 (75.42) 62.08 (59.10)
The results reveal several interesting patterns: (1) MMLU and MMLU-Pro features show reasonable cross-transferability, likely due to their shared multiple-choice format and reasoning requirements, (2) BBQ features demonstrate good transferability to MMLU tasks, suggesting that bias mitigation features capture general reasoning capabilities, and (3) features optimized for specific tasks consistently outperform transferred features, validating the importance of task-specific feature selection. These findings support our discussion of limited but meaningful transferability among structurally similar tasks.
A.6 Text Classification Validation
To validate the effectiveness of correlation-based feature selection, we conduct controlled experiments on text classification tasks where ground truth labels provide clear supervision signals. The experiments utilize GPT-2 (Radford et al., 2019) with publicly available SAEs from Bloom et al. (Bloom, 2024) on the bias-focused text classification dataset EMGSD (King et al., 2024).
For each bias category, we extract the most correlated features using max-pooling over all text tokens, then apply steering by either adding positively correlated features or subtracting negatively correlated features. Steering effectiveness is evaluated using the same classifier employed in the original dataset.
Table 9: Bias steering effectiveness across different demographic categories on EMGSD dataset. Mitigation reduces bias scores, while amplification increases them.
Mitigation Amplification
Category Baseline CorrSteer Biased CorrSteer
Gender 0.177 0.616 0.897 0.922
LGBTQ+ 0.091 0.561 0.941 0.882
Nationality 0.125 0.732 0.937 0.945
Profession 0.128 0.625 0.890 0.921
Race 0.308 0.769 0.846 0.846
Religion 0.109 0.655 0.945 0.928
Results demonstrate that correlation-selected features provide effective steering control across all demographic categories (Table 9). Our steering approach effectively reduces bias, with steered outputs showing substantially lower bias scores compared to biased baselines, supporting the generalizability of our approach across different domains and SAE training paradigms.
A demonstration of our bias mitigation results is available at https://huggingface.co/spaces/seonglae/CorrSteer, showcasing real-time steering capabilities.
A.7 Complete Feature Lists
This section presents the complete feature lists for each task, showing the top-1 features aggregated from all layers. Each feature is labeled with the format L{layer}/{index} to identify its layer and index position. Features selected by CorrSteer-P after pruning are highlighted in bold.
Each feature entry includes the feature description along with its coefficient and correlation value. SAE feature descriptions are obtained through the Neuronpedia API (https://www.neuronpedia.org/), providing automated semantic interpretations of selected features. Feature indices are hyperlinked to their corresponding Neuronpedia pages for detailed analysis.
Feature descriptions that are well-aligned with the target task are highlighted in bold, and the highest correlations for each task are also emphasized in bold. Following each layer’s highest correlated feature, we include additional relevant features listed below.
A.7.1 Gemma-2B
BBQ (Ambiguous)
Figure: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of Gemma 2 2B.
  • L1/6088 specific formatting or structural elements within text, such as timestamps and code (coeff: 2.280, corr: 0.134)
  • L2/15089 key actions and processes related to achievements and collaboration (coeff: 4.898, corr: 0.166)
  • L3/6151 references to statistical or numerical data in research contexts (coeff: 3.537, corr: 0.091)
  • L4/11047 certain types of mathematical or programming syntax (coeff: 2.854, corr: 0.121)
  • L5/7502 expressions of honesty and self-awareness in discourse (coeff: 3.117, corr: 0.199)
  • L6/324 structured sentences that present facts, warnings, or errors, often with an emphasis on important details (coeff: 2.886, corr: 0.169)
  • L7/4487 the presence of detailed structured elements within a document, such as headings or separators in a legal or formal layout (coeff: 4.996, corr: 0.102)
  • L8/4669 special tokens or specific formatting in the text (coeff: 4.378, corr: 0.147)
  • L9/1435 elements related to copyright and licensing information (coeff: 8.737, corr: 0.107)
  • L10/4557 interactions involving guessing or determining the correctness of information (coeff: 4.246, corr: 0.202)
  • L11/6144 return statements in code (coeff: 4.347, corr: 0.192)
  • L12/15862 punctuation marks and formatting elements in the text (coeff: 2.718, corr: 0.214)
  • L13/4379 punctuation symbols and their frequency (coeff: 6.779, corr: 0.165)
  • L14/12922 dialogue or conversational exchanges involving questioning and responses (coeff: 1.754, corr: 0.181)
  • L15/12813 medical terms related to respiratory health and conditions (coeff: 3.537, corr: 0.242)
  • L16/9006 declarations regarding conflicts of interest and funding in research publications (coeff: 2.606, corr: 0.330)
  • L17/11021 phrases related to scientific research and findings (coeff: 6.777, corr: 0.554)
  • L18/14447 references to medical data and statistics (coeff: 9.667, corr: 0.533)
  • L19/11289 assignment and return statements in programming contexts (coeff: 10.429, corr: 0.538)
  • L20/2040 occurrences of logical values and conditions in programming or data handling contexts (coeff: 9.166, corr: 0.523)
  • L21/8433 keywords related to programming functions and their definitions (coeff: 5.983, corr: 0.440)
  • L22/10377 code snippets that include assignments and return statements (coeff: 14.919, corr: 0.458)
  • L23/6394 structured data or code-like formats (coeff: 34.482, corr: 0.442)
  • L24/14051 references to education systems and their impact on health initiatives (coeff: 25.098, corr: 0.413)
  • L25/12534 references to emotional states or descriptions of personal experiences (coeff: 18.414, corr: 0.394)
Additional relevant features:
  • L8/8123 questions that ask for truthfulness or correctness regarding options or statements (coeff: 3.725, corr: -0.133)
  • L17/9134 choice-related phrases and expressions of preference (coeff: 2.379, corr: -0.451)
  • L19/15745 phrases related to decision-making and choice, particularly in the context of parenting and social interactions (coeff: 9.740, corr: -0.464)
Figure: Top correlated features with BBQ ambig on frequency in each layer of Gemma 2 2B.
BBQ (Disambiguous)
Figure: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of Gemma 2 2B.
  • L1/7001 code structure and elements in programming, particularly related to class and variable definitions (coeff: 2.126, corr: 0.114)
  • L2/8432 HTML and JavaScript code related to the Bootstrap framework (coeff: 2.418, corr: 0.140)
  • L3/10179 terms related to health and medical supplements (coeff: 2.383, corr: 0.134)
  • L4/3444 various types of headers, specifically those that denote responses and results within the context of exchanges or interactions (coeff: 2.192, corr: 0.114)
  • L5/697 terms related to price dynamics and economic relationships (coeff: 3.766, corr: 0.088)
  • L6/2491 references to sources or citations in a document (coeff: 2.618, corr: 0.110)
  • L7/6269 references to visual elements such as figures and tables (coeff: 1.293, corr: 0.135)
  • L8/5927 mathematical examples and notations (coeff: 3.347, corr: 0.259)
  • L9/7854 structures related to the declaration and manipulation of result variables in a programming context (coeff: 10.475, corr: 0.189)
  • L10/15705 references to file operations and data management in code (coeff: 6.145, corr: 0.215)
  • L11/13926 mathematical expressions and calculations (coeff: 8.203, corr: 0.154)
  • L12/1085 references to court cases and legal statutes (coeff: 1.839, corr: 0.220)
  • L13/536 technical details related to manufacturing processes (coeff: 4.417, corr: 0.178)
  • L14/10612 structured data or code snippets related to databases (coeff: 5.030, corr: 0.225)
  • L15/2822 structured data formats or code snippets related to programming (coeff: 1.632, corr: 0.176)
  • L16/6602 the presence of specific numerical or coding patterns in data (coeff: 6.773, corr: 0.300)
  • L17/5137 mathematical symbols and functions related to field theories (coeff: 8.483, corr: 0.559)
  • L18/3178 code or programming-related elements (coeff: 7.851, corr: 0.507)
  • L19/11641 technical components or elements in code (coeff: 16.336, corr: 0.414)
  • L20/12748 structured data representations and their attributes (coeff: 28.025, corr: 0.394)
  • L21/14337 code-related keywords and method definitions in programming contexts (coeff: 20.453, corr: 0.392)
  • L22/13921 elements related to database structure and definitions (coeff: 18.510, corr: 0.420)
  • L23/12349 technical terms related to software or code management (coeff: 5.893, corr: 0.331)
  • L24/16355 definitions and mathematical notation in text (coeff: 39.910, corr: 0.326)
  • L25/4307 occurrences of programming syntax related to object-oriented structures (coeff: 19.460, corr: 0.384)
Additional relevant features:
  • L18/1127 references to gender and associated options/choices in forms (coeff: 4.813, corr: 0.207)
  • L19/15745 phrases related to decision-making and choice, particularly in the context of parenting and social interactions (coeff: 11.875, corr: 0.226)
  • L23/12048 terms related to racism and social injustice (coeff: 2.661, corr: 0.147)
Figure: Top correlated features with BBQ disambig on frequency in each layer of Gemma 2 2B.
HarmBench
Figure: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of Gemma 2 2B.
  • L1/9572 occurrences of the semicolon character (coeff: 5.206, corr: 0.692)
  • L2/6712 references to worship and its related symbols or icons (coeff: 5.699, corr: 0.692)
  • L3/16207 syntax elements and formatting in code or mathematical expressions (coeff: 2.583, corr: 0.686)
  • L4/3109 forms of the verb "to be" and its variations (coeff: 5.891, corr: 0.696)
  • L5/11099 sentences that include personal affirmations or declarations of identity (coeff: 16.934, corr: 0.737)
  • L6/12241 instances of the verb "to be" in various forms and their contexts (coeff: 7.338, corr: 0.735)
  • L7/11722 phrases related to legal terms and the rejection of arguments in court cases (coeff: 5.035, corr: 0.779)
  • L8/8642 expressions of self-identity and subjective experience (coeff: 8.729, corr: 0.745)
  • L9/9298 strongly negative or dismissive opinions about claims and arguments (coeff: 7.525, corr: 0.775)
  • L10/3037 references to legal issues and compliance (coeff: 6.667, corr: 0.723)
  • L11/6905 statements of identity and self-description (coeff: 13.810, corr: 0.735)
  • L12/12039 phrases related to providing assistance and support (coeff: 5.253, corr: 0.741)
  • L13/6715 text that discusses accountability and the need for forgiveness (coeff: 6.992, corr: 0.709)
  • L14/2949 statements and phrases related to political criticism and condemnation (coeff: 16.620, corr: 0.739)
  • L15/1570 judgments regarding moral and ethical standards related to exploitation and human rights issues (coeff: 23.824, corr: 0.742)
  • L16/5113 expressions of personal identity and emotional states (coeff: 21.832, corr: 0.743)
  • L17/5887 references to tools and functional capabilities related to programming or software development (coeff: 11.389, corr: 0.720)
  • L18/1411 negative statements or denials (coeff: 20.537, corr: 0.712)
  • L19/324 phrases related to legal procedures and considerations (coeff: 35.610, corr: 0.710)
  • L20/5192 questions that seek clarification or challenge assumptions (coeff: 45.662, corr: 0.718)
  • L21/7129 negative sentiments and expressions of doubt or denial (coeff: 33.225, corr: 0.721)
  • L22/3311 references to food and culinary experiences (coeff: 19.000, corr: 0.746)
  • L23/11246 instances of strong negative sentiment or rejection (coeff: 61.642, corr: 0.711)
  • L24/12773 first-person pronouns and references to personal experiences or actions (coeff: 50.332, corr: 0.699)
  • L25/3912 negative sentiments or refusals (coeff: 57.431, corr: 0.711)
Figure: Top correlated features with HarmBench on frequency in each layer of Gemma 2 2B.
MMLU
Figure: Top correlated features with selected features from CorrSteer-P with MMLU on coefficient in each layer of Gemma 2 2B.
  • L1/13714 colons and semicolons used in lists or programming syntax (coeff: 0.403, corr: 0.140)
  • L2/6273 specific medical terminology and its implications (coeff: 1.548, corr: 0.175)
  • L3/12378 programming-related elements and commands (coeff: 1.094, corr: 0.164)
  • L4/11047 certain types of mathematical or programming syntax (coeff: 2.944, corr: 0.225)
  • L5/8581 phrases that indicate research findings or results (coeff: 0.077, corr: 0.115)
  • L6/5275 sentences expressing doubt or conditionality in arguments (coeff: 4.939, corr: 0.140)
  • L7/14726 periods and other punctuation marks that signify sentence endings or significant separations in text (coeff: 2.532, corr: 0.159)
  • L8/15039 terms related to research methodologies and experimental design (coeff: 0.309, corr: 0.152)
  • L9/15654 variations of the word "correct" in various contexts (coeff: 0.414, corr: 0.136)
  • L10/11729 coding attributes and properties related to light types in a 3D programming context (coeff: 2.919, corr: 0.174)
  • L11/13204 code syntax and structure, particularly related to variable assignments and function calls (coeff: 5.369, corr: 0.126)
  • L12/6392 XML-like structured data elements (coeff: 1.033, corr: 0.200)
  • L13/12281 mathematical expressions and concepts related to positive values (coeff: 0.919, corr: 0.254)
  • L14/7 significant scientific findings and their specific details (coeff: 6.002, corr: 0.170)
  • L15/8678 phrases related to announcements or updates (coeff: 4.906, corr: 0.281)
  • L16/12421 programming constructs and their structures within code snippets (coeff: 5.593, corr: 0.251)
  • L17/13214 error messages and diagnostic codes (coeff: 9.790, corr: 0.294)
  • L18/1127 references to gender and associated options/choices in forms (coeff: 4.805, corr: 0.376)
  • L19/2174 input fields and value assignments in a form-like structure (coeff: 8.405, corr: 0.402)
  • L20/12748 structured data representations and their attributes (coeff: 20.884, corr: 0.394)
  • L21/14337 code-related keywords and method definitions in programming contexts (coeff: 13.228, corr: 0.362)
  • L22/5939 technical jargon and terminology related to chemistry and biochemistry (coeff: 5.582, corr: 0.313)
  • L23/10424 statistical terms and symbols related to data analysis and significance testing (coeff: 25.724, corr: 0.400)
  • L24/16355 definitions and mathematical notation in text (coeff: 36.077, corr: 0.367)
  • L25/10388 phrases related to health-related actions and topics (coeff: 33.899, corr: 0.336)
Figure: Top correlated features with MMLU on frequency in each layer of Gemma 2 2B.
MMLU-Pro
Figure: Top correlated features with selected features from CorrSteer-P with MMLU-Pro on coefficient in each layer of Gemma 2 2B.
  • L1/9317 phrases related to changes in social and organizational dynamics (coeff: 1.859, corr: 0.169)
  • L2/3714 mathematical notation, specifically related to set notation and expressions involving functions (coeff: 0.761, corr: 0.226)
  • L3/11980 statements providing answers or conclusions regarding questions or hypotheses (coeff: 3.699, corr: 0.153)
  • L4/15960 terms related to medical procedures and conditions (coeff: 6.817, corr: 0.170)
  • L5/7502 expressions of honesty and self-awareness in discourse (coeff: 2.187, corr: 0.086)
  • L6/6201 numeric representations of system specifications or configurations (coeff: 14.877, corr: 0.210)
  • L7/8790 structured data formats and their attributes (coeff: 1.209, corr: 0.182)
  • L8/11297 structured data and programming constructs (coeff: 2.176, corr: 0.209)
  • L9/15336 references to mathematical or computational problems and their solutions (coeff: 6.407, corr: 0.200)
  • L10/10805 terms related to medical conditions and biological factors (coeff: 1.277, corr: 0.237)
  • L11/1909 affirmative or negative responses in the context of questions (coeff: 2.296, corr: 0.226)
  • L12/14752 legal and governmental terms related to authority and judgment (coeff: 1.369, corr: 0.253)
  • L13/12991 mathematical operations and expressions (coeff: 2.560, corr: 0.239)
  • L14/10780 comments and documentation markers in code (coeff: 1.455, corr: 0.252)
  • L15/2262 references to variable declarations and data structures in programming contexts (coeff: 1.183, corr: 0.334)
  • L16/3142 mathematical symbols and notation used in equations (coeff: 5.691, corr: 0.285)
  • L17/1175 mathematical expressions and applications related to programming or data structures (coeff: 3.091, corr: 0.483)
  • L18/682 function declarations and their return types in a programming context (coeff: 3.406, corr: 0.448)
  • L19/11641 technical components or elements in code (coeff: 2.144, corr: 0.414)
  • L20/12748 structured data representations and their attributes (coeff: 7.134, corr: 0.529)
  • L21/1944 code structures and syntax related to programming and mathematics (coeff: 9.251, corr: 0.456)
  • L22/12947 scientific terminology related to healthcare and medical research (coeff: 11.241, corr: 0.440)
  • L23/5752 associations and relationships among scientific variables and observations (coeff: 10.133, corr: 0.497)
  • L24/8188 syntax related to code structure and operations (coeff: 11.861, corr: 0.482)
  • L25/8643 scientific terms and concepts related to biochemistry and cellular processes (coeff: 11.439, corr: 0.545)
Figure: Top correlated features with MMLU-Pro on frequency in each layer of Gemma 2 2B.
GSM8K
Figure: Top correlated features with selected features from CorrSteer-P with GSM8K on coefficient in each layer of Gemma 2 2B.
  • L1/13475 specific quantitative or statistical information (coeff: 9.936, corr: 0.251)
  • L2/2098 references to leadership and management isolation in workplace contexts (coeff: 3.080, corr: 0.180)
  • L3/8338 significant quantities within code snippets, likely indicating important operations or constructs (coeff: 6.302, corr: 0.250)
  • L4/687 HTML tags and attributes related to layout and styling (coeff: 2.037, corr: 0.188)
  • L5/697 terms related to price dynamics and economic relationships (coeff: 6.091, corr: 0.193)
  • L6/13460 references to safety and regulatory issues in automobile contexts (coeff: 9.501, corr: 0.219)
  • L7/9514 structured data or code snippets, potentially relating to geographical regions and associated identifiers (coeff: 1.309, corr: 0.167)
  • L8/2024 names of notable performance venues and cultural institutions (coeff: 14.384, corr: 0.210)
  • L9/15115 discussions related to crime scene investigations and forensic evidence (coeff: 5.074, corr: 0.188)
  • L10/2794 elements of conversation or dialogue (coeff: 5.602, corr: 0.188)
  • L11/7313 mathematical equations and expressions (coeff: 26.252, corr: 0.176)
  • L12/12707 technical or scientific terminology related to systems and processes (coeff: 2.860, corr: 0.245)
  • L13/14319 code snippets and their associated structures within documents (coeff: 2.731, corr: 0.253)
  • L14/4217 expressions of emotional reactions and feedback (coeff: 3.772, corr: 0.246)
  • L15/1685 instances of structured data or messages indicating communication or queries (coeff: 7.282, corr: 0.255)
  • L16/14919 instances of unique identifiers or markers in a dataset (coeff: 24.774, corr: 0.223)
  • L17/7185 curly braces and structured programming syntax elements (coeff: 6.245, corr: 0.252)
  • L18/3732 code syntax elements such as brackets and semicolons (coeff: 4.064, corr: 0.249)
  • L19/2015 structures related to function definitions and method calls in programming code (coeff: 8.802, corr: 0.277)
  • L20/15616 elements of code structure and syntax in programming contexts (coeff: 4.350, corr: 0.258)
  • L21/12547 phrases and words that express confusion or dissatisfaction with situations (coeff: 24.211, corr: 0.251)
  • L22/7903 mathematical notation and symbols used in equations (coeff: 7.295, corr: 0.313)
  • L23/12425 mathematical expressions and symbols (coeff: 19.202, corr: 0.294)
  • L24/2274 programming syntax and structure specific to coding languages (coeff: 10.205, corr: 0.348)
  • L25/3469 technical aspects related to semiconductor devices and their manufacturing processes (coeff: 23.158, corr: 0.284)
Figure: Top correlated features with GSM8K on frequency in each layer of Gemma 2 2B.
SimpleQA
Figure: Top correlated features with selected features from CorrSteer-P with SimpleQA on coefficient in each layer of Gemma 2 2B.
  • L1/14904 references to Congress and legislative processes (coeff: 0.263, corr: 0.192)
  • L2/1089 terms and concepts related to integrals and the importance of integration in various contexts (coeff: 0.225, corr: 0.228)
  • L3/12843 terms related to durability and long-lasting qualities (coeff: 0.219, corr: 0.178)
  • L4/680 references to standards, particularly in legal and medical contexts (coeff: 0.055, corr: 0.194)
  • L5/4460 references to legal cases and court rulings (coeff: 0.093, corr: 0.193)
  • L6/9777 expressions of agreement or dissent and the context surrounding them (coeff: 0.153, corr: 0.192)
  • L7/2431 phrases related to time management and constraints (coeff: 0.253, corr: 0.213)
  • L8/14209 HTML coding elements and formatting commands (coeff: 0.097, corr: 0.194)
  • L9/7856 terms related to penalties and scoring in sporting events (coeff: 0.257, corr: 0.196)
  • L10/2446 terms related to health and legal matters (coeff: 0.350, corr: 0.177)
  • L11/7954 legal terminology and references to court cases and proceedings (coeff: 0.164, corr: 0.194)
  • L12/1495 phrases related to the duration and continuity of experiences over time (coeff: 0.208, corr: 0.195)
  • L13/12119 references to various parameters and aspects within scientific or technical contexts (coeff: 0.148, corr: 0.194)
  • L14/6355 references to India and its cultural context (coeff: 0.377, corr: 0.201)
  • L15/4385 programming constructs related to class and method declarations in Java (coeff: 0.177, corr: 0.194)
  • L16/7182 expressions of enthusiasm or amazement (coeff: 0.457, corr: 0.216)
  • L17/6346 references to offices or organizational structures (coeff: 0.274, corr: 0.194)
  • L18/5258 terms related to legal charges and prosecutions (coeff: 0.843, corr: 0.200)
  • L19/4202 technical terms and concepts related to physical phenomena and their mathematical descriptions (coeff: 0.262, corr: 0.194)
  • L20/6557 legal terms and phrases related to criminal charges and legal proceedings (coeff: 0.953, corr: 0.220)
  • L21/13830 references to political leaders and government roles (coeff: 8.090, corr: 0.196)
  • L22/15897 phrases related to international relations and cooperation, particularly in the context of political statements and actions (coeff: 0.648, corr: 0.237)
  • L23/15190 terms related to health, well-being, and interventions for obesity and mental illness (coeff: 0.832, corr: 0.237)
  • L24/15228 references to political figures and their actions (coeff: 2.043, corr: 0.222)
  • L25/2531 expressions of political opinion regarding government spending and fiscal policies (coeff: 1.664, corr: 0.200)
Figure: Top correlated features with SimpleQA on frequency in each layer of Gemma 2 2B.
XSTest
Figure: Top correlated features with selected features from CorrSteer-P with XSTest on coefficient in each layer of Gemma 2 2B.
  • L1/4509 terms and concepts related to scientific and mathematical structures and functions (coeff: 1.940, corr: 0.333)
  • L2/4679 financial metrics and forecasts related to stock performance (coeff: 1.584, corr: 0.301)
  • L3/1326 legal terminology and references to statutes and claims (coeff: 4.088, corr: 0.256)
  • L4/12152 references to geographical locations and their associated attributes (coeff: 0.961, corr: 0.259)
  • L5/5939 terms related to signals and their coding in biological contexts (coeff: 3.391, corr: 0.248)
  • L6/4376 numerical values and specific formatting related to data structures or coding (coeff: 6.713, corr: 0.259)
  • L7/4886 representations of numerical data, particularly in scientific contexts (coeff: 3.074, corr: 0.286)
  • L8/10825 punctuation marks and special characters (coeff: 5.194, corr: 0.296)
  • L9/9228 punctuation marks, especially periods and quotation marks (coeff: 4.712, corr: 0.323)
  • L10/13244 information related to military casualties and incidents (coeff: 2.760, corr: 0.270)
  • L11/5734 sections or punctuation that denote lists or explanations (coeff: 4.304, corr: 0.243)
  • L12/12342 symbols and mathematical notation related to expressions or equations in mathematical contexts (coeff: 15.373, corr: 0.282)
  • L13/10964 mathematical terms and symbols (coeff: 16.622, corr: 0.274)
  • L14/7655 structured data, such as XML or JSON formats (coeff: 16.195, corr: 0.275)
  • L15/5114 terms related to evaluation and validation processes (coeff: 23.117, corr: 0.248)
  • L16/1547 code or programming-related syntax (coeff: 21.527, corr: 0.283)
  • L17/10813 references to movies, actors, and significant film industry terms (coeff: 9.662, corr: 0.243)
  • L18/8615 legal terminology and concepts related to judicial authority and precedent (coeff: 9.006, corr: 0.282)
  • L19/2998 elements related to research findings, including factors, conclusions, and reasoning (coeff: 13.956, corr: 0.245)
  • L20/9419 names of individuals and titles (coeff: 10.648, corr: 0.272)
  • L21/15170 isolated segments of code or technical content (coeff: 36.804, corr: 0.264)
  • L22/11042 punctuation marks that indicate the start or end of lists or key points in a text (coeff: 28.482, corr: 0.294)
  • L23/8993 structured API documentation elements and syntax (coeff: 23.447, corr: 0.280)
  • L24/4448 terms related to scientific analysis and results reporting (coeff: 16.649, corr: 0.287)
  • L25/7968 elements related to health assessments and metrics (coeff: 9.863, corr: 0.307)
Figure: Top correlated features with XSTest on frequency in each layer of Gemma 2 2B.
A.7.2 Llama-3.1-8B
BBQ (Ambiguous)
Figure: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of Llama 3.1 8B.
  • L1/23207 phrases related to legal or regulatory frameworks (coeff: 0.463, corr: 0.111)
  • L2/2680 titles and key information related to television series episodes (coeff: 0.002, corr: 0.117)
  • L3/23846 discussions around societal structures and issues related to mental health and crime (coeff: 0.487, corr: 0.127)
  • L4/30896 occurrences of numerical values and references to measurements (coeff: 0.089, corr: 0.128)
  • L5/18555 instances of past and present tense verbs, particularly focusing on actions and conditions (coeff: 0.193, corr: 0.137)
  • L6/25246 technical terms and code snippets related to software development and programming logic (coeff: 0.277, corr: 0.147)
  • L7/11878 specific numerical identifiers and related metadata in technical documents (coeff: 0.365, corr: 0.178)
  • L8/4790 keywords related to data structures and programming concepts (coeff: 0.172, corr: 0.163)
  • L9/2700 references to extraterrestrial or paranormal beings and phenomena (coeff: 0.354, corr: 0.187)
  • L10/23355 phrases or constructs that emphasize comparison or simile (coeff: 0.812, corr: 0.168)
  • L11/18132 references to specific books, movies, or artworks (coeff: 0.167, corr: 0.181)
  • L12/14096 references to specific locations or settings in various contexts (coeff: 0.084, corr: 0.189)
  • L13/26526 references to error handling in programming (coeff: 0.493, corr: 0.203)
  • L14/13393 statistical percentages and survey data (coeff: 0.192, corr: 0.324)
  • L15/25166 themes of neutrality and balance in discourse (coeff: 0.259, corr: 0.433)
  • L16/21816 phrases related to financial or economic assessments (coeff: 0.543, corr: 0.363)
  • L17/5782 references to equality and equity in rights and opportunities (coeff: 0.368, corr: 0.298)
  • L18/28196 references to knowledge, learning, and understanding in various contexts (coeff: 0.303, corr: 0.390)
  • L19/29460 discussions about extremes and balance (coeff: 0.811, corr: 0.440)
  • L20/13319 expressions of mixed opinions or complex character evaluations (coeff: 1.413, corr: 0.473)
  • L21/8518 references to articles and citations in academic databases (coeff: 2.719, corr: 0.349)
  • L22/28263 percentages and statistical data concerning opinions or responses (coeff: 1.024, corr: 0.464)
  • L23/638 formal structures and procedures within organizational contexts (coeff: 1.054, corr: 0.496)
  • L24/19174 code constructs and control flow keywords related to conditions and returns (coeff: 1.890, corr: 0.465)
  • L25/10753 expressions of perception or belief in social dynamics (coeff: 1.147, corr: 0.428)
  • L26/27899 code structure and logical operations involving object hierarchy and data types (coeff: 1.025, corr: 0.452)
  • L27/1765 quantitative data related to project development and financial metrics (coeff: 2.597, corr: 0.384)
  • L28/21019 financial data and statistics related to development projects (coeff: 0.856, corr: 0.323)
  • L29/17998 code snippets related to JavaScript or Java programming functions and structures (coeff: 1.735, corr: 0.385)
  • L30/17084 numerical data related to financial projections and resource development (coeff: 1.308, corr: 0.390)
  • L31/10728 auxiliary verbs and words indicating obligation or possibility (coeff: 1.530, corr: 0.239)
Figure: Top correlated features with BBQ ambig on frequency in each layer of Llama 3.1 8B.
BBQ (Disambiguous)
Figure: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of Llama 3.1 8B.
  • L1/5891 technical terms and references in programming and development contexts (coeff: 0.154, corr: 0.086)
  • L2/21865 references to essays, articles, and related writing concepts (coeff: 0.784, corr: 0.084)
  • L3/3413 elements related to user engagement and user-friendly design (coeff: 0.332, corr: 0.100)
  • L4/3712 elements related to programming and computation (coeff: 0.458, corr: 0.086)
  • L5/18066 references to educational administration and school district issues (coeff: 0.229, corr: 0.118)
  • L6/28294 references to machine learning models and recommendation systems (coeff: 0.301, corr: 0.119)
  • L7/7762 specific language constructs related to coordination and organization (coeff: 0.416, corr: 0.124)
  • L8/25466 terms related to hierarchical structures or classifications (coeff: 1.032, corr: 0.124)
  • L9/5313 key concepts related to project management and planning (coeff: 0.645, corr: 0.139)
  • L10/13407 negative actions and attitudes that hinder interpersonal relationships and community engagement (coeff: 0.256, corr: 0.152)
  • L11/18350 references to institutions and systems regarding public services (coeff: 0.900, corr: 0.128)
  • L12/13336 phrases and concepts related to community and social interactions (coeff: 0.377, corr: 0.144)
  • L13/15793 negation phrases and words indicating absence or lack (coeff: 0.695, corr: 0.167)
  • L14/31962 details related to physical displacement or movement in a spatial context (coeff: 1.384, corr: 0.217)
  • L15/2128 references to programming elements and constructs (coeff: 0.977, corr: 0.277)
  • L16/6219 code-related syntax and structures within programming languages (coeff: 0.830, corr: 0.292)
  • L17/12610 technical terminology related to programming and software development (coeff: 0.706, corr: 0.275)
  • L18/16458 HTML tags and structured data elements (coeff: 2.113, corr: 0.285)
  • L19/6432 numerical values and the structure of dates or game scores (coeff: 0.909, corr: 0.284)
  • L20/28406 tokens related to timestamps, specifically date and time formats (coeff: 0.942, corr: 0.297)
  • L21/15538 references to time management techniques and motivational strategies (coeff: 0.388, corr: 0.199)
  • L22/11286 monetary amounts or financial figures (coeff: 0.531, corr: 0.245)
  • L23/30672 phrases involving the concept of answers or responses (coeff: 1.211, corr: 0.222)
  • L24/5888 references to answers or responses in discussions or questions (coeff: 1.152, corr: 0.222)
  • L25/22713 mathematical notations and symbols (coeff: 1.235, corr: 0.253)
  • L26/22133 names of authors and their affiliations in academic contexts (coeff: 1.953, corr: 0.269)
  • L27/12321 structural elements and parameters in programming code or data structures (coeff: 0.539, corr: 0.180)
  • L28/23202 specific numbers and their context within factual statements (coeff: 1.897, corr: 0.267)
  • L29/3168 keywords related to health and medical terminology (coeff: 3.175, corr: 0.253)
  • L30/22450 terms and phrases related to health and medical conditions (coeff: 3.219, corr: 0.167)
  • L31/18173 procedural commands and technical instructions related to software and settings (coeff: 1.440, corr: 0.188)
Figure: Top correlated features with BBQ disambig on frequency in each layer of Llama 3.1 8B.
HarmBench
Figure: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of Llama 3.1 8B.
  • L1/15747 repetitive phrases or expressions related to certainty or emphasis (coeff: 0.491, corr: 0.524)
  • L2/25715 references to collective experiences and communal responsibility (coeff: 1.032, corr: 0.590)
  • L3/23621 negations and assertions related to existence and actions (coeff: 1.116, corr: 0.580)
  • L4/26750 first-person pronouns indicating personal experiences and thoughts (coeff: 3.468, corr: 0.586)
  • L5/300 instances of political criticism and hypocrisy (coeff: 1.587, corr: 0.734)
  • L6/21616 discussions about legality, morality, and the implications of actions in ethical contexts (coeff: 1.458, corr: 0.590)
  • L7/17622 phrases related to trust and loyalty in political contexts (coeff: 1.128, corr: 0.639)
  • L8/6508 expressions related to the condemnation of sexual assault and violence (coeff: 1.322, corr: 0.648)
  • L9/27026 concepts related to limits and responsibilities in relationships and societal interactions (coeff: 1.425, corr: 0.619)
  • L10/9364 expressions of moral outrage and condemnation regarding social and ethical issues (coeff: 1.324, corr: 0.633)
  • L11/16561 expressions of personal opinion and moral judgments (coeff: 1.810, corr: 0.608)
  • L12/5839 strong statements against violence and discrimination (coeff: 1.271, corr: 0.694)
  • L13/15443 emotional expressions of affection or attachment (coeff: 1.637, corr: 0.569)
  • L14/22046 phrases and sentiments associated with moral judgments and emotional responses (coeff: 0.750, corr: 0.582)
  • L15/5498 phrases related to environmental and climate impact (coeff: 0.696, corr: 0.609)
  • L16/8375 topics related to stigma and mental health awareness (coeff: 0.938, corr: 0.614)
  • L17/15876 expressions of self-doubt or uncertainty (coeff: 0.582, corr: 0.660)
  • L18/6210 phrases related to educational support and challenges faced by teachers (coeff: 0.964, corr: 0.641)
  • L19/5854 references to seeking medical advice and guidance (coeff: 1.148, corr: 0.564)
  • L20/11388 elements related to moral and ethical dilemmas (coeff: 3.490, corr: 0.633)
  • L21/9674 references to racism and social justice issues (coeff: 0.712, corr: 0.559)
  • L22/4650 expressions of self-awareness and personal growth mixed with skepticism towards collective beliefs (coeff: 2.235, corr: 0.560)
  • L23/28291 phrases discussing social justice and advocacy for marginalized communities (coeff: 2.165, corr: 0.636)
  • L24/21055 phrases related to self-identity and personal reflection (coeff: 2.357, corr: 0.679)
  • L25/16450 themes of emotional struggle and interpersonal relationships (coeff: 2.415, corr: 0.602)
  • L26/6648 phrases indicating moral judgment or hypocrisy in political discourse (coeff: 1.541, corr: 0.593)
  • L27/10654 expressions of emotional conflict and personal reflection (coeff: 1.653, corr: 0.655)
  • L28/522 themes of courage and resilience in writing (coeff: 0.915, corr: 0.578)
  • L29/13883 complex emotional responses and reflections on interpersonal relationships (coeff: 2.977, corr: 0.639)
  • L30/4588 expressions of emotional needs and desires in relationships (coeff: 1.480, corr: 0.586)
  • L31/31181 references to familial relationships and memorial details (coeff: 1.218, corr: 0.639)
Figure: Top correlated features with HarmBench on frequency in each layer of Llama 3.1 8B.
MMLU
Figure: Top correlated features with selected features from CorrSteer-P with MMLU ambig on coefficient in each layer of Llama 3.1 8B.
  • L1/4557 specific numeric values and measurements related to instructions or guidelines (coeff: 0.695, corr: 0.094)
  • L2/27893 terms related to technology, specifically graphics processing units (GPUs) and their applications (coeff: 0.348, corr: 0.157)
  • L3/204 terms and concepts related to financial metrics and performance evaluation (coeff: 1.037, corr: 0.139)
  • L4/23545 questions that lead to detailed inquiries or clarifications (coeff: 1.142, corr: 0.131)
  • L5/17458 terms related to theoretical concepts and methodologies in scientific discussions (coeff: 0.497, corr: 0.124)
  • L6/650 specific identifiers, particularly those related to content or lists (coeff: 0.780, corr: 0.110)
  • L7/13659 references to lists, particularly those pertaining to security or classification contexts (coeff: 0.885, corr: 0.118)
  • L8/1649 key terms related to organizational assistance and functionality within various contexts (coeff: 0.871, corr: 0.116)
  • L9/19730 various forms of interviews and discussions related to current events or cultural topics (coeff: 0.397, corr: 0.108)
  • L10/20495 terms related to requirements and definitions within various contexts (coeff: 0.949, corr: 0.099)
  • L11/20851 legal and academic terminology related to charges and reports (coeff: 0.897, corr: 0.100)
  • L12/26346 specific nouns and proper names related to various contexts (coeff: 0.454, corr: 0.104)
  • L13/551 terms related to medical results and actions taken toward health management (coeff: 0.830, corr: 0.143)
  • L14/11013 phrases indicating relationships between people or entities (coeff: 0.366, corr: 0.165)
  • L15/9446 expressions of passion and enthusiasm in various contexts (coeff: 0.327, corr: 0.195)
  • L16/6219 code-related syntax and structures within programming languages (coeff: 1.094, corr: 0.274)
  • L17/26604 references to programming concepts and structures (coeff: 0.957, corr: 0.301)
  • L18/28750 structured data elements and patterns, possibly related to programming or data analysis (coeff: 0.936, corr: 0.288)
  • L19/6432 numerical values and the structure of dates or game scores (coeff: 1.587, corr: 0.365)
  • L20/28406 tokens related to timestamps, specifically date and time formats (coeff: 1.051, corr: 0.319)
  • L21/15538 references to time management techniques and motivational strategies (coeff: 1.014, corr: 0.347)
  • L22/11286 monetary amounts or financial figures (coeff: 1.269, corr: 0.322)
  • L23/15096 phrases related to significant life events and milestones (coeff: 1.125, corr: 0.281)
  • L24/18010 references to dates and significant life events (coeff: 1.631, corr: 0.256)
  • L25/22713 mathematical notations and symbols (coeff: 1.209, corr: 0.287)
  • L26/22133 names of authors and their affiliations in academic contexts (coeff: 2.331, corr: 0.331)
  • L27/19268 references to academic qualifications, research, and involvement in educational activities (coeff: 0.826, corr: 0.310)
  • L28/23202 specific numbers and their context within factual statements (coeff: 2.318, corr: 0.307)
  • L29/3168 keywords related to health and medical terminology (coeff: 3.545, corr: 0.255)
  • L30/23403 terms associated with uncertainty and error (coeff: 0.986, corr: 0.274)
  • L31/6722 instances of code-related syntax and formatting (coeff: 0.538, corr: 0.159)
Figure: Top correlated features with MMLU on frequency in each layer of Llama 3.1 8B.
MMLU-Pro
Figure: Top correlated features with selected features from CorrSteer-P with MMLU-Pro ambig on coefficient in each layer of Llama 3.1 8B.
  • L1/2403 specific numeric values and measurements related to instructions or guidelines (coeff: 0.286, corr: 0.216)
  • L2/85 phrases related to service expectations and quality assurance
    (coeff: 0.212, corr: 0.259)
  • L3/204 terms and concepts related to financial metrics and performance evaluation (coeff: 0.996, corr: 0.265)
  • L4/14539 content related to sources and references in articles (coeff: 0.432, corr: 0.250)
  • L5/2831 references to urgency and scheduling events (coeff: 0.348, corr: 0.277)
  • L6/7784 instances of various relational and transactional terms within context (coeff: 0.153, corr: 0.265)
  • L7/22238 references to examples or lists in discussions or reports (coeff: 0.446, corr: 0.282)
  • L8/7704 keywords related to television series and their reception
    (coeff: 0.630, corr: 0.244)
  • L9/4007 references to various types of businesses and their classifications (coeff: 0.298, corr: 0.248)
  • L10/3783 key phrases and concepts related to business development and investment processes (coeff: 0.454, corr: 0.281)
  • L11/7301 components of structured data or content organization (coeff: 0.807, corr: 0.261)
  • L12/28750 financial terms and conditions related to trading or commerce (coeff: 0.563, corr: 0.306)
  • L13/16587 phrases indicating action or involvement in events or developments (coeff: 0.366, corr: 0.285)
  • L14/28135 references to specific geographic locations or entities (coeff: 0.490, corr: 0.312)
  • L15/9446 expressions of passion and enthusiasm in various contexts (coeff: 0.425, corr: 0.337)
  • L16/6219 code-related syntax and structures within programming languages (coeff: 0.342, corr: 0.323)
  • L17/26604references to programming concepts and structures (coeff: 0.469, corr: 0.357)
  • L18/2624 references to criminal activity and associated legal consequences (coeff: 0.478, corr: 0.371)
  • L19/6432 numerical values and the structure of dates or game scores (coeff: 0.966, corr: 0.381)
  • L20/28406 tokens related to timestamps, specifically date and time formats (coeff: 0.628, corr: 0.368)
  • L21/15538 references to time management techniques and motivational strategies (coeff: 0.391, corr: 0.345)
  • L22/11286 monetary amounts or financial figures (coeff: 0.697, corr: 0.380)
  • L23/21146 programming and coding structures, particularly related to network protocols and data handling (coeff: 0.853, corr: 0.348)
  • L24/7967 references to specific locations or addresses (coeff: 0.837, corr: 0.350)
  • L25/16619 instances of authorship and attribution in the text (coeff: 0.864, corr: 0.347)
  • L26/22133 names of authors and their affiliations in academic contexts(coeff: 0.813, corr: 0.413)
  • L27/19268 references to academic qualifications, research, and involvement in educational activities (coeff: 0.318, corr: 0.271)
  • L28/23202 specific numbers and their context within factual statements (coeff: 1.120, corr: 0.304)
  • L29/12442 patterns related to digital platforms and software updates (coeff: 2.528, corr: 0.249)
  • L30/19427 specific numerical values and statistical data (coeff: 0.374, corr: 0.311)
  • L31/9926 numbers, particularly in relation to financial data and statistics (coeff: 10.348, corr: 0.280)
Figure: Top correlated features with MMLU-Pro on frequency in each layer of Llama 3.1 8B.
SimpleQA
Figure: Top correlated features with SimpleQA on frequency in each layer of Llama 3.1 8B.
  • L1/28160 references to height, specifically focusing on the term "tall" (coeff: 1.580, corr: 0.454)
  • L2/16190 references to geographical locations, particularly islands
    (coeff: 0.148, corr: 0.383)
  • L3/24193 references to deserts and desert-related imagery (coeff: 0.541, corr: 0.496)
  • L4/25100 references to dumpster rental services and pricing (coeff: 0.205, corr: 0.457)
  • L5/15924 the occurrence of the word "in" and its context within the text (coeff: 0.396, corr: 0.418)
  • L6/7008 references to artificial entities and technologies (coeff: 2.402, corr: 0.383)
  • L7/6257 terms and phrases related to artificial elements or creations (coeff: 2.049, corr: 0.381)
  • L8/30264 phrases or terms that indicate suitability or excellence in context (coeff: 0.029, corr: 0.377)
  • L9/23784 programming-related keywords and constructs (coeff: 0.089, corr: 0.377)
  • L10/30120 phrases that encourage action or reminders related to specific tasks (coeff: 0.057, corr: 0.377)
  • L11/962 conjunctions that introduce reasoning or causation (coeff: 0.396, corr: 0.410)
  • L12/31391 references to authors and their written works (coeff: 0.472, corr: 0.437)
  • L13/19013 references to biological family classifications (coeff: 2.618, corr: 0.387)
  • L14/12579 references to global outreach and international presence (coeff: 0.077, corr: 0.377)
  • L15/18867 references to biological classifications, specifically family names in taxonomy (coeff: 2.004, corr: 0.386)
  • L16/22032 biological classifications of species, particularly family and genus names (coeff: 2.364, corr: 0.417)
  • L17/30566 phrases related to ownership or affiliation (coeff: 0.884, corr: 0.377)
  • L18/24624 specific terms associated with the media and entertainment industry (coeff: 0.952, corr: 0.410)
  • L19/25841 references to personal growth and transformation experiences (coeff: 1.140, corr: 0.395)
  • L20/23840 references to legislative districts and redistricting processes (coeff: 0.438, corr: 0.409)
  • L21/9851 references to volcanic activity (coeff: 0.258, corr: 0.377)
  • L22/20579 references to educational programs and initiatives (coeff: 0.744, corr: 0.400)
  • L23/11708 complex arguments and perspectives in academic discourse (coeff: 0.323, corr: 0.423)
  • L24/14877 specific procedural or data-related elements in formal documents (coeff: 0.292, corr: 0.530)
  • L25/18055 words associated with appreciation and commendation (coeff: 0.542, corr: 0.469)
  • L26/10617 emotional expressions and relationships in personal narratives (coeff: 0.317, corr: 0.435)
  • L27/135 activities related to travel and tourism (coeff: 0.924, corr: 0.380)
  • L28/29877 references to the concept of "home." (coeff: 0.964, corr: 0.377)
  • L29/4392 references to clothing and dress codes, particularly in relation to gender identity and expression (coeff: 0.410, corr: 0.382)
  • L30/22633 public methods in a programming context (coeff: 0.310, corr: 0.377)
  • L31/6171 references to artificial intelligence and its related concepts (coeff: 1.429, corr: 0.377)
Figure: Top correlated features with SimpleQA on frequency in each layer of Llama 3.1 8B.
XSTest
Figure: Top correlated features with XSTest on frequency in each layer of Llama 3.1 8B.
  • L1/6754 references to studies and publications (coeff: 0.256, corr: 0.367)
  • L2/5332 names and characteristics associated with aviation or flight (coeff: 0.276, corr: 0.331)
  • L3/16461 terms related to marine life and conservation efforts (coeff: 1.265, corr: 0.394)
  • L4/2446 proper nouns and specific entities (coeff: 0.310, corr: 0.334)
  • L5/25000 names of notable individuals and places related to historical or cultural significance (coeff: 0.862, corr: 0.354)
  • L6/10424 information related to personal details and statistics about individuals (coeff: 0.220, corr: 0.355)
  • L7/20235 words and phrases associated with measurement or assessment (coeff: 0.784, corr: 0.364)
  • L8/22807 concepts related to capital budgeting and investment decision-making (coeff: 0.420, corr: 0.411)
  • L9/16423 references to specific organizations, laws, or conditions related to societal issues (coeff: 0.636, corr: 0.455)
  • L10/11238 phrases related to collaboration and community involvement (coeff: 0.880, corr: 0.365)
  • L11/29172 legal terminology related to civil rights and obligations (coeff: 0.618, corr: 0.383)
  • L12/19663 negative descriptors or concepts related to cowardice and existence (coeff: 0.735, corr: 0.384)
  • L13/19506 numeric or alphanumeric strings and specific identifiers (coeff: 0.608, corr: 0.403)
  • L14/13505 structured question-answer formats and indicators of a discussion or inquiry (coeff: 4.659, corr: 0.369)
  • L15/23853 references to female characters and their relationships in narratives (coeff: 0.682, corr: 0.400)
  • L16/1652 names and identifiers related to locations and organizations (coeff: 1.220, corr: 0.373)
  • L17/21476 references to influential figures in scientific history and significant concepts from their work (coeff: 2.046, corr: 0.357)
  • L18/25543 names and specific references related to individuals, locations, and organizations in a political context (coeff: 0.941, corr: 0.353)
  • L19/2102 significant historical events and their impact on society (coeff: 1.691, corr: 0.366)
  • L20/21486 various references to awards, accolades, and notable achievements within literary and cinematic contexts (coeff: 2.183, corr: 0.385)
  • L21/8477 references to influential figures and their contributions in various contexts (coeff: 2.008, corr: 0.383)
  • L22/16870 references to disasters and their impacts (coeff: 2.837, corr: 0.366)
  • L23/15524 references to specific events or characters in films (coeff: 1.834, corr: 0.400)
  • L24/15231 references to specific events or characters in films (coeff: 1.747, corr: 0.392)
  • L25/16855 references to corporate entities and financial transactions (coeff: 0.763, corr: 0.375)
  • L26/1578 references to specific individuals or organizations involved in social causes or environmental conservation (coeff: 0.948, corr: 0.338)
  • L27/11758 connections to authoritative figures and organizational roles (coeff: 1.300, corr: 0.367)
  • L28/425 instances of specific names and organizational references in a text (coeff: 2.291, corr: 0.360)
  • L29/17372 terms related to health and illness (coeff: 0.888, corr: 0.312)
  • L30/11223 titles and descriptors of programs or services related to community support (coeff: 4.643, corr: 0.352)
  • L31/2111 descriptions and features of software products (coeff: 1.614, corr: 0.276)
Figure: Top correlated features with XSTest on frequency in each layer of Llama 3.1 8B.

### v2

@@
-CorrSteer: Steering Improves Task Performance and Safety in LLMs through Correlation-based Sparse Autoencoder Feature Selection
+CorrSteer: Generation-Time LLM Steering via Correlated Sparse Autoencoder Features
 Abstract
@@
 To address these limitations, we propose CorrSteer, which selects features by correlating sample correctness with SAE activations from generated tokens at inference time.
-This approach uses only inference-time activations to extract more relevant features, thereby avoiding spurious correlations.
+This approach uses only inference-time activations to extract more relevant features, thereby reducing spurious correlations.
 It also obtains steering coefficients from average activations, automating the entire pipeline.
-Our method shows improved task performance on QA, bias mitigation, jailbreaking prevention, and reasoning benchmarks on Gemma 2 2B and LLaMA 3.1 8B, notably achieving a +4.1% improvement in MMLU performance and a +22.9% improvement in HarmBench with only 4000 samples.
+Our method shows improved task performance on QA, bias mitigation, jailbreaking prevention, and reasoning benchmarks on Gemma-2 2B and LLaMA-3.1 8B, notably achieving a +3.3% improvement in MMLU performance with 4000 samples and a +27.2% improvement in HarmBench with only 108 samples.
 Selected features demonstrate semantically meaningful patterns aligned with each task’s requirements, revealing the underlying capabilities that drive performance.
 Our work establishes correlation-based selection as an effective and scalable approach for automated SAE steering across language model applications.
-CorrSteer: Steering Improves Task Performance and Safety in LLMs through Correlation-based Sparse Autoencoder Feature Selection
-Seonglae Cho1,2  Zekun Wu1,2  Adriano Koshiyama1,2 1Holistic AI  2University College London
+Figure 1: System diagram of CorrSteer.
+CorrSteer selects task-relevant SAE features by correlating generated-token activations with outcomes, and constructs steering vectors applied as CorrSteer-S, CorrSteer-A, or CorrSteer-P.
+This generation-time steering shifts model behavior from unintended to intended responses while reducing side effects.
 1 Introduction
 Sparse Autoencoders (SAEs) have emerged as a powerful tool for decomposing superposed representations in large language models (LLMs) into interpretable sparse latent dimensions (Huben et al., 2023).
-By reconstructing neural activations through a sparse bottleneck, SAEs effectively disentangle semantic features that can be leveraged for downstream tasks such as probing and steering (Bricken et al., 2023).
+By reconstructing neural activations through a sparse bottleneck, SAEs disentangle semantic features that can be leveraged for downstream tasks such as probing and steering (Bricken et al., 2023).
 However, existing SAE-based steering approaches face significant limitations: (1) contrastive datasets (Soo et al., 2025) or large activation storage (Zhao et al., 2025; Arad et al., 2025) are required to identify the direction of the steering, and (2) they rely on the hidden states of context tokens to select both the features and their coefficients.
-Figure 1: Top correlated features with MMLU in each layer of Gemma 2 2B.
-Consequently, current use cases of SAE-based steering have been restricted to specific applications, such as bias mitigation (Durmus et al., 2024), knowledge unlearning (Muhamed et al., 2025; Wang et al., 2025; Zhou et al., 2025; Cywiński and Deja, 2025), and jailbreaking prevention (O’Brien et al., 2025).
+Consequently, current use cases of SAE-based steering have been restricted to specific applications, such as bias mitigation (Durmus et al., 2024), knowledge unlearning (Muhamed et al., 2025; Wang et al., 2025; Zhou et al., 2025; Cywiński & Deja, 2025), and jailbreaking prevention (O’Brien et al., 2025).
 Moreover, SAE feature selection in these applications does not directly reflect language models’ generation capabilities, potentially limiting their applicability.
-Figure 2: Benchmark performance of CorrSteer variants compared with the baseline on Gemma 2 2B.
 To address these limitations, this work introduces CorrSteer, which leverages generation-time features by correlating with task outcomes for task-specific feature selection and steering coefficient determination.
@@
 Finally, by defining SER (Side Effect Ratio), three variants of CorrSteer are compared targeting the minimization of SER against fine-tuning.
-2 Background
+2 Related Work
 Mechanistic Interpretability aims to reverse-engineer neural networks into human-interpretable components (Olah et al., 2020; Elhage et al., 2021).
@@
 This efficient representation strategy complicates efforts to identify the consistent role of specific latent dimensions.
-2.1 Sparse Autoencoders
-Sparse Autoencoders (Huben et al., 2023; Bricken et al., 2023) address the superposition problem by learning to decompose neural activations into interpretable, sparse features.
-Given an activation vector $\mathbf{x}\in\mathbb{R}^{d}$, an SAE learns an encoder $f_{\text{enc}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ and decoder $f_{\text{dec}}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ where $k\gg d$, such that:
-$\mathbf{z}$ $=f_{\text{enc}}(\mathbf{x})=\text{Activation}(\mathbf{W}_{\text{enc}}\mathbf{x}+\mathbf{b}_{\text{enc}})$ (1)
-$\hat{\mathbf{x}}$ $=f_{\text{dec}}(\mathbf{z})=\mathbf{W}_{\text{dec}}\mathbf{z}+\mathbf{b}_{\text{dec}}$ (2)
-The training objective is usually a combination of reconstruction loss with sparsity regularization:
-$\mathcal{L}=\|\mathbf{x}-\hat{\mathbf{x}}\|^{2}+\lambda\|\mathbf{z}\|_{1}$ (3)
-2.2 Steering Vectors
 Steering Vectors (Subramani et al., 2022) represent a class of methods for controlling neural network outputs by manipulating internal activations.
 Traditional approaches, such as CAA (Rimsky et al., 2024), compute activation differences between contrasting examples and apply these differences.
-For precise control to not inadvertently affect related behaviors simultaneously, PaCE (Luo et al., 2024) utilizes sparse coding for orthogonal steering directions.
-2.3 SAE-based Steering
-SAE-based steering leverages SAE latents for predictable control based on feature semantics.
+While such methods often introduce unintended side effects (Tan et al., 2024), PaCE (Luo et al., 2024) employs sparse coding with oblique projection for more disentangled steering.
+SAE-based Steering leverages Sparse Autoencoder latents for predictable control based on feature semantics.
 SAE-TS (Chalnev et al., 2024; Soo et al., 2025) reduces the side effects of steering by linearly approximating feature directions.
@@
 Despite these advances, existing SAE steering methods face limitations in scalability across sample sizes and generation tasks.
-3 Method
-Linear correlation offers both interpretability and faithfulness as a criterion for feature selection.
-SAEs capture linear relationships, consistent with the Linear Representation Hypothesis (Socher et al., 2013; Faruqui et al., 2015; Park et al., 2023), and have a proven ability to disentangle interpretable features in a linear manner.
-The faithfulness of Pearson correlation is further supported by recent work from Oikarinen et al.
-(2025).
-3.1 Correlation-based Feature Selection
-Our approach, CorrSteer, centers on the observation that features most correlated with task performance are likely to be relevant for steering.
-The approach employs the Pearson correlation coefficient, applied only to generation-time features—specifically to the last token at each step.
-Given a set of SAE features $\mathbf{z}=[z_{1},z_{2},\ldots,z_{D}]$ and corresponding task performance scores $\mathbf{y}=[y_{1},y_{2},\ldots,y_{n}]$ for $n$ samples, the correlation for each feature $i$ is computed as:
+Recent work has shown that SAEs capture linear relationships consistent with the Linear Representation Hypothesis (Socher et al., 2013; Faruqui et al., 2015; Park et al., 2023), and Pearson correlation has been demonstrated as a faithful measure for such linear dependencies (Oikarinen et al., 2025).
+These findings motivate our proposed approach, CorrSteer, which leverages correlation-based feature selection for automated and scalable SAE steering.
+This simplicity, combined with scalability and interpretability, distinguishes CorrSteer from prior SAE steering methods
+3 The CorrSteer Method
+Figure 1 provides an overview of CorrSteer, illustrating how correlation-based feature selection and steering interventions are applied.
+CorrSteer is a simple yet scalable pipeline that steers language models by linking generation-time SAE activations with task outcomes.
+Our method first identifies task-relevant features via correlation, then assigns coefficients from their natural activation scales, and finally applies steering vectors during inference.
+This design emphasizes three advantages over prior SAE-based steering: simplicity, scalability, and interpretability.
+3.1 Correlation-Guided Feature Selection
+The central idea of CorrSteer is that features most correlated with task performance are also the most promising candidates for steering.
+Pearson correlation is well-suited for SAE’s inherently linear architecture where features are designed to be linearly combined (Bricken et al., 2023), aligning with the Linear Representation Hypothesis (Park et al., 2023) and leveraging correlation as a faithful measure for linear dependencies in neural representations (Oikarinen et al., 2025).
+To capture this relationship, we compute correlations only on generation-time activations, focusing on the last generated token at each step, since these activations are most directly tied to model output correctness.
+Formally, given a set of SAE features $\mathbf{z}=[z_{1},z_{2},\ldots,z_{D}]$ and corresponding task performance scores $\mathbf{y}=[y_{1},y_{2},\ldots,y_{n}]$ for $n$ samples, the correlation for each feature $i$ is computed as:
 $r_{i}=\frac{\text{Cov}(z_{i},y)}{\sqrt{\text{Var}(z_{i})\cdot\text{Var}(y)}}$ (1)
-To handle the computational challenges of large SAE feature dictionaries (typically $10^{4}$$10^{5}$ features), a streaming correlation accumulator is implemented that maintains $O(1)$ memory complexity:
+To handle the computational challenges of large SAE feature dictionaries (typically $10^{4}$-$10^{5}$ features), a streaming correlation accumulator is implemented that maintains $O(1)$ memory complexity:
 Algorithm 1 Streaming Correlation Computation
- Initialize: $\sum x_{i}=0,\sum x_{i}^{2}=0,\sum x_{i}y_{i}=0,\sum y_{i}=0,\sum y_{i}^{2}=0,n=0$
+ Initialize accumulators:
+$\sum x_{i}=0,\;\;\sum x_{i}^{2}=0,\;\;\sum x_{i}y_{i}=0,\;\;\sum y_{i}=0,\;\;\sum y_{i}^{2}=0,\;\;n=0$
 for each batch $(\mathbf{X}_{\text{batch}},\mathbf{y}_{\text{batch}})$ do
   Update running sums for each feature dimension
-  $n\leftarrow n+\text{batch\_size}$
+  $n\leftarrow n+|\mathbf{y}_{\text{batch}}|$
 end for
- Compute correlations:
-$medmath{r_{i}=\frac{n\sum x_{i}y_{i}-\sum x_{i}\sum y_{i}}{\sqrt{(n\sum x_{i}^{2}-(\sum x_{i})^{2})(n\sum y_{i}^{2}-(\sum y_{i})^{2})}}}$
-This computation is $O(1)$ with respect to sample size, and $O(LD)$ for fixed layer count $L$ and SAE latent dimension $D$.
+ Compute correlations for each feature $i$:
+$r_{i}=\frac{n\sum x_{i}y_{i}-\sum x_{i}\sum y_{i}}{\sqrt{(n\sum x_{i}^{2}-(\sum x_{i})^{2})(n\sum y_{i}^{2}-(\sum y_{i})^{2})}}$
+This computation maintains $O(1)$ space complexity with respect to sample size, while time complexity is $O(N)$ for $N$ samples, and $O(LD)$ for fixed layer count $L$ and SAE latent dimension $D$.
 For generation tasks requiring multiple tokens, max-pooling is employed over valid token positions to aggregate feature activations, as empirically validated in our pooling comparison study (Table 5).
-3.2 Coefficient Calculation
-The steering coefficient for a selected feature $i$ is computed as the mean activation value among samples where task performance is positive.
-This approach is preferable to contrastive-based calculation, since SAE features produce non-negative activations due to their ReLU-based activation functions (Bricken et al., 2023) and thus cannot be meaningfully subtracted in a contrastive manner; negative activations are often unrelated noise (Joseph Bloom, 2024).
-$c_{i}=\frac{1}{|\{j:y_{j}>0\}|}\sum_{j:y_{j}>0}z_{i,j}$ (5)
-This ensures that the steering magnitude reflects the natural activation scale of the feature during successful task performance.
-3.3 Steering Implementation
-During inference, steering is applied by modifying residual stream activations.
-For a selected feature $i$, with coefficient $c_{i}$ and SAE decoder weights $\mathbf{W}_{\text{dec}}$ (the feature direction (Templeton et al., 2024)), the steering vector is:
-$\mathbf{v}_{\text{steer}}=c_{i}\cdot\mathbf{W}_{\text{dec}}[:,i]$ (6)
-The modified activation is:
-$\mathbf{x}^{\prime}=\mathbf{x}+\mathbf{v}_{\text{steer}}$ (7)
-Steering is applied to tokens corresponding to the positions from which the features were originally extracted, rather than only to the last token (Luo et al., 2024; Rimsky et al., 2024) or every token (Soo et al., 2025).
-3.4 Feature Extraction Strategies
-Figure 3: Comparison of selected features between CorrSteer-1, CorrSteer-A, and CorrSteer-P on the BBQ disambiguous task across all layers of Gemma 2 2B.
-Red points indicate features selected by each method.
-For each layer $\ell$, we obtain SAE activations from the residual stream and rank features by correlation with task performance.
-The method compares a global view aggregated across layers and a layer-wise view for selecting features to steer.
-Three strategies are implemented:
-  • CorrSteer-1: Select the single highest-correlated feature from the global view, allowing cross-layer feature competition.
-  • CorrSteer-A: Select the top-1 feature within each layer.
-  • CorrSteer-P: Prune features from CorrSteer-A using a validation set, retaining only those that improve baseline performance to reduce side effects from spurious correlations.
-Each method has distinct pros and cons, discussed in Section 5.
-Only positively correlated features are selected to ensure steering induces positive activation, as our ablation study confirms that negative correlation features consistently degrade performance 6.
-All methods are fully automated based on observed activations without hyperparameter tuning.
-3.5 Side Effect Ratio
-A key challenge in correlation-based feature selection is distinguishing features that causally contribute to task success from those that merely correlate due to the model’s internal state, potentially causing unintended side effects.
-To quantify side effects, the Side Effect Ratio (SER) is defined as the proportion of negatively changed answers among all changed answers:
-$SER=\frac{\text{\# negatively changed answers}}{\text{\# all changed answers}}$ (8)
-This measure does not isolate the side effect of each individual feature; rather, it serves as a combined metric reflecting how well the selected features are optimized for the task without degrading the model’s original abilities.
-To reduce side effects, the approach focuses on features activated during the generation process, under the hypothesis that generation-time activations are more likely to be causally relevant to output.
+3.2 Coefficient Estimation from Positive Outcomes
+For each selected feature $i$, we define its steering coefficient as the mean activation over samples with positive task outcomes. Formally:
+$c_{i}=\frac{1}{|\{j:y_{j}>0\}|}\sum_{j:y_{j}>0}z_{i,j}.$ (2)
+This formulation directly anchors the steering magnitude to the feature’s natural activation scale during successful performance.
+Unlike contrastive-based methods, it leverages the non-negativity of SAE activations (arising from ReLU) (Bricken et al., 2023), thereby avoiding ill-posed subtraction between activation states and ensuring stable, semantically faithful steering.
+These coefficients are then used at inference time to construct steering vectors that modify the model’s residual stream.
+3.3 Inference-Time Steering Mechanism
+At inference time, steering modifies residual stream activations at each token generation step.
+For a selected feature $i$ with coefficient $c_{i}$ and SAE decoder weights $\mathbf{W}_{\text{dec}}$ (its feature direction (Templeton et al., 2024)), the steering vector $\mathbf{v}_{\text{steer}}=c_{i}\cdot\mathbf{W}_{\text{dec}}[:,i]$ is added to the residual stream as $\mathbf{x}^{\prime}=\mathbf{x}+\mathbf{v}_{\text{steer}}$ for each generated token position.
+We apply steering exclusively at token positions where the corresponding features were extracted, rather than uniformly across all tokens (Soo et al., 2025) or restricted to the final token (Luo et al., 2024; Rimsky et al., 2024).
+Since many benchmarks involve multi-token generations, this raises the question of how to aggregate activations across tokens when computing correlations and coefficients, which we address next.
+3.4 Pooling Strategy for Feature Aggregation.
+Two pooling strategies are available for coefficient and correlation calculations: max-pooling and mean-pooling.
+For multi-token generation tasks, max-pooling consistently outperforms mean-pooling, as empirically demonstrated in Table 5, likely due to its better capture of peak feature activations relevant to task success.
+However, for coefficient calculation in longer generation tasks such as GSM8K reasoning, mean-pooling is preferred as max-pooling produces excessively large coefficient values.
+Applying these large coefficients to every generated token degrades performance, leading to the adoption of mean-pooling for reasoning tasks.
+Having established how features are aggregated across tokens, we next turn to how features are selected across layers.
+3.5 Automated Multi-Layer Feature Selection
+For each layer $\ell$, we extract SAE activations from the residual stream and rank features by their correlation with task performance.
+We consider both a global view aggregating correlations across layers and a layer-wise view that preserves layer-specific structure.
+Based on these perspectives, we implement three fully automated strategies (no hyperparameter tuning required):
+  • CorrSteer-S. Select the single most positively correlated feature across all layers (global view). This minimal variant tests whether a single feature suffices for causal performance improvements.
+  • CorrSteer-A. Select the top positively correlated feature from each layer. This design probes whether layer-wise features collectively form circuits that enhance task performance.
+  • CorrSteer-P. Begin with CorrSteer-A and apply validation-based pruning, retaining only those features that improve over the non-steered model. This enables finer-grained subcircuit analysis.
+Only positively correlated features are retained, as ablation experiments confirm that negatively correlated features consistently degrade performance (Appendix 6).
+Figure 2 illustrates these strategies on the BBQ (disambiguous) task across all layers of Gemma-2 2B, highlighting how CorrSteer-S, CorrSteer-A, and CorrSteer-P differ in terms of selected feature distribution (red points).
+While CorrSteer-S focuses on a single dominant signal, CorrSteer-A distributes selections across layers, and CorrSteer-P prunes this set to retain only features that yield improvements.
+These differences highlight distinct trade-offs in global versus layer-wise selection.
+However, feature selection may also introduce unintended side effects, which we address next.
+Figure 2: Comparison of features selected by CorrSteer-S, CorrSteer-A, and CorrSteer-P on BBQ (disambiguous) across all Gemma-2 2B layers. Red points denote selected features.
+3.6 Quantifying Side Effects via SER
+Correlation-based feature selection risks capturing spurious associations rather than causal drivers, leading to unintended degradations. We quantify this with the Side Effect Ratio (SER), defined as
+$\text{SER}=\frac{\#\text{ negatively changed answers}}{\#\text{ all changed answers}}.$ (3)
+Lower SER indicates more reliable steering, isolating features that improve performance without harmful side effects.
+This measure does not isolate the side effect of each individual feature; rather, it serves as a combined metric reflecting how well selected features are optimized for the task without degrading the model’s original abilities.
+To reduce side effects, the approach focuses on features activated during generation, under the hypothesis that generation-time activations are more likely causally relevant to output.
 This inference-time focus is empirically validated by our pooling experiments (Table 5).
-Additionally, in the multi-layer approach, a validation-based filtering mechanism is introduced (CorrSteer-P), retaining only features that demonstrate actual steering effectiveness.
-4 Experiments
-Figure 4: Relation between consumed sample counts and test performance, final matched count of the selected features, and the most correlated features from each layer of Gemma 2 2B.
-The dotted lines show the baseline of default LLM performance and constrained decoding performance on MMLU answer options.
-4.1 Experimental Setup
-CorrSteer is evaluated across diverse generation benchmarks to demonstrate practical effectiveness.
-Models and SAEs: Experiments are conducted using Gemma-2 2B (Team, 2024a) and LLaMA-3.1 8B (Team, 2024b) models, paired with their corresponding SAE releases from Gemma Scope (Lieberum et al., 2024) and LLaMA Scope (He et al., 2024), respectively.
+Additionally, in the multi-layer approach, a validation-based filtering mechanism is introduced (CorrSteer-P), retaining only features that demonstrate steering effectiveness.
+4 Experimental Setup
+Experiments are conducted using Gemma-2 2B (Team, 2024a) and LLaMA-3.1 8B (Team, 2024b) models, paired with their corresponding SAE releases from Gemma Scope (Lieberum et al., 2024) and LLaMA Scope (He et al., 2024), respectively.
 Both SAE families employ JumpReLU activation (Rajamanoharan et al., 2024).
-Additionally, the Gemma-2-2B-IT model with SAEs is employed, leveraging the fact that SAEs are typically transferable across fine-tuned models (Kissane et al., 2024), with proven low loss reported in the Gemma Scope paper (Lieberum et al., 2024).
-Datasets: Our evaluation encompasses multiple benchmark categories:
-  • Question Answering: MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024)
-  • Reasoning: GSM8k (Cobbe et al., 2021)
-  • Bias Evaluation: BBQ (Parrish et al., 2022)
-  • Factuality: SimpleQA (Wei et al., 2024)
-  • Safety: HarmBench (Mazeika et al., 2024), XSTest (Röttger et al., 2024)
-For safety benchmarks, both the refusal benchmark HarmBench and the overrefusal benchmark XSTest are included to evaluate not only the steering ability for rejection but also the model’s capability to identify the context of requests.
-Evaluation Metrics: For multiple-choice tasks (MMLU, MMLU-Pro, BBQ), exact match accuracy is used.
+Additionally, the Gemma-2 2B-IT model with SAEs is employed, leveraging the fact that SAEs are typically transferable across fine-tuned models (Kissane et al., 2024), with proven low loss reported in the Gemma Scope paper (Lieberum et al., 2024).
+Evaluation Benchmarks We evaluate CorrSteer on a suite of benchmarks spanning five categories:
+  • Knowledge: MMLU (Hendrycks et al., 2021) and MMLU-Pro (Wang et al., 2024) test broad-domain expertise under zero-shot settings.
+  • Reasoning: GSM8K (Cobbe et al., 2021) probes multi-step mathematical reasoning ability.
+  • Bias: BBQ (Parrish et al., 2022) measures sensitivity to social bias and stereotypes.
+  • Factuality: SimpleQA (Wei et al., 2024) assesses short-form factual consistency.
+  • Safety: HarmBench (Mazeika et al., 2024) and XSTest (Röttger et al., 2024) evaluate resistance to unsafe or sensitive content generation.
+For safety benchmarks, both HarmBench (refusal) and XSTest (overrefusal) evaluate steering ability and contextual understanding.
+Side Effect Evaluation.
+We measure Side Effect Ratio (SER) to quantify unintended performance degradations (Table 2).
+CorrSteer’s SER is compared against fine-tuning baselines across question-answering datasets.
+Additionally, we validate our positive-only feature selection by comparing performance when using negatively correlated features (Appendix 6).
+We also assess different pooling strategies to verify that inference-time token selection is optimal (Table 5).
+Pooling Strategies for Feature Aggregation.
+To verify that our pooling design in Section 3.4 is robust, we conduct an ablation comparing three strategies for aggregating SAE activations across tokens: (i) mean-pooling, which averages activations across tokens; (ii) all-token pooling, which aggregates contributions from every position; and (iii) max-pooling, which selects the strongest activation.
+We evaluate these alternatives on GSM8K (reasoning), BBQ (bias), and HarmBench/XSTest (safety), covering both single-token and multi-token generation tasks.
+This setup isolates the effect of pooling and allows us to test whether CorrSteer’s empirically motivated default choices are consistently optimal across task types.
+Feature Interpretability and Transferability Analysis.
+Performance-improving features are analyzed post-hoc using Neuronpedia descriptions to examine whether correlation-selected features exhibit semantic coherence (Appendix A.8.1).
+We analyze whether performance-improving features correspond to meaningful behaviors such as refusal, neutrality, or structured reasoning.
+Safe/unsafe tendency inspection and task-wise breakdowns test whether CorrSteer activates task-relevant semantics rather than spurious signals.
+Finally, we probe transferability by evaluating features selected on one benchmark (e.g., MMLU) on others (e.g., BBQ, MMLU-Pro) to test whether our method identifies generalizable circuits (Table 8).
+5 Results and Discussion
+Table 1 and Table 3 present comprehensive results across evaluation benchmarks. CorrSteer demonstrates improvements across question answering, bias mitigation, and safety benchmarks.
+Table 1: Performance comparison across CorrSteer variants and other steering methods on Gemma-2 2B.
+Results are reported as mean ± standard deviation across 5 random seeds (3 for GSM8K).
+Within each method category, the best results are highlighted in bold, and the second-best results are highlighted in italics.
+Method MMLU MMLU-Pro SimpleQA BBQ Ambig BBQ Disambig HarmBench XSTest GSM8K
+CorrSteer Variants
+Non-steered 52.21 ± 0.04 30.40 ± 0.21 3.78 ± 0.17 59.46 ± 0.21 75.38 ± 0.14 46.61 ± 2.78 86.35 ± 0.32 54.44 ± 0.35
+CorrSteer-S 52.99 ± 0.47 30.38 ± 0.08 3.68 ± 0.07 62.39 ± 0.02 75.70 ± 0.01 46.61 ± 0.76 86.77 ± 0.48 53.63 ± 0.72
+CorrSteer-P 54.70 ± 1.22 30.63 ± 0.13 3.80 ± 0.14 66.00 ± 2.15 76.48 ± 0.64 66.08 ± 20.20 86.46 ± 0.37 53.10 ± 0.74
+CorrSteer-A 55.48 ± 0.59 30.93 ± 0.19 3.74 ± 0.07 62.06 ± 0.84 76.53 ± 0.23 73.75 ± 8.84 86.98 ± 1.45 40.34 ± 24.43
+Other Methods
+Fine-tuning 55.75 ± 0.09 35.32 ± 2.70 47.00 ± 0.33
+SPARE (MI) 54.97 ± 0.87 30.84 ± 0.18 3.72 ± 0.04 64.81 ± 2.12 76.25 ± 0.59 65.43 ± 14.34 86.82 ± 0.76
+DSG (Fisher) 52.81 ± 0.59 30.33 ± 0.16 3.66 ± 0.06 61.75 ± 1.39 75.61 ± 0.16 45.86 ± 1.76 86.35 ± 0.59
+CAA 55.13 ± 1.00 28.01 ± 5.79 3.71 ± 0.07 62.40 ± 1.07 76.32 ± 0.40 43.14 ± 28.95 72.95 ± 17.50
+5.1 Comparison with Baselines
+Across benchmarks, CorrSteer-A and CorrSteer-P achieve the strongest results, with CorrSteer-P showing particular dominance in LLaMA-3.1 8B.
+This can be attributed to the less disentangled nature of LLaMA Scope features under superposition, which necessitates more aggressive pruning.
+Results on both Gemma-2 2B and LLaMA-3.1 8B confirm consistent improvement patterns.
+The correlation-based approach consistently outperforms mutual information (MI) and Fisher information-based methods, supporting the faithfulness of SAE’s linear representation.
+This suggests that linear correlation-based feature extraction aligns with the linear latent space of SAEs, where features are designed to be linearly combined.
+Existing steering approaches rely on contrastive examples restricted to static contexts, while CorrSteer directly leveraging generation-time activations, extending SAE-based steering and achieving practical gains across QA, safety, and bias benchmarks.
+Head-to-head comparison with CAA (Rimsky et al., 2024), DSG (Muhamed et al., 2025), or SPARE (Zhao et al., 2025) is not directly applicable since these methods require contrastive datasets rather than generation-time features.
+However, for comparison purposes, we apply our generation-time feature selection approach to these methods.
+For fair comparison, we applied the same test-time features and average positive coefficients across methods, with MI and Fisher information-based methods using substituted feature selection while CAA directly uses correct and incorrect answer activation differences.
+Furthermore, other methods also show improved performance when adapted to use generation-time features, demonstrating the effectiveness of our generation-time feature selection approach independent of the specific steering mechanism.
+While fine-tuning achieves higher raw accuracy, CorrSteer offers advantages in side-effect reduction.
+On MMLU, CorrSteer-A achieves competitive accuracy (55.48% vs.
+55.75%) while halving SER (0.20 vs.
+0.41) (Table 1, Table 2).
+Although fine-tuning outperforms CorrSteer variants in raw accuracy on GSM8K and MMLU-Pro, CorrSteer maintains substantially lower SER across tasks.
+Moreover, CorrSteer can be layered on top of fine-tuned models as complementary enhancement.
+Feature Collaboration and Circuit Effects.
+CorrSteer-A demonstrates superior performance in 5 out of 8 tasks, indicating that improvements often emerge from feature collaboration within circuits, even when individual feature steering yields limited benefit.
+Multi-layer approaches such as CorrSteer-A and CorrSteer-P consistently outperform the single-layer CorrSteer-S, aligning with prior findings on circuit-level interventions (Liu et al., 2024; Zhao et al., 2025).
+Safety and Factuality.
+On HarmBench, selected features enhance refusal ability, achieving a 27.2% gain, though this primarily reflects increased refusal rather than fine-grained safety.
+In contrast, XSTest shows limited gains due to the benchmark’s over-refusal bias.
+This outcome is expected given the static nature of CorrSteer, which cannot easily separate benign from harmful requests.
+Similarly, on SimpleQA, CorrSteer yields only marginal improvement, confirming that the method enhances adherence to task requirements without introducing external factual knowledge.
+This is desirable, as it suggests CorrSteer modifies behavior rather than injecting content absent from the base model.
+5.2 Efficiency and Scalability
+CorrSteer serves as an auxiliary mechanism that identifies task-relevant features through generation-time correlations, complementing supervised fine-tuning and remaining effective when applied on top of fine-tuned models.
+The pipeline is fully automated, requires no hyperparameter tuning, and generalizes across tasks and domains with minimal adjustment.
+The streaming correlation algorithm operates with $\mathcal{O}(1)$ memory complexity relative to dataset size, ensuring scalability to large corpora.
+CorrSteer performs effectively with as few as 100 samples, though stable performance requires approximately 4,000 samples (Appendix A.1).
+Once steering vectors are extracted, inference requires no SAE dependency, since fixed feature sets and coefficients are sufficient.
+Figure 3: Relation between sample counts and test performance, final matched count of selected features, and most correlated features from each Gemma-2 2B layer.
+Dotted lines show baseline default LLM performance and constrained decoding performance on MMLU answer options.
+Training Sample Requirements:
+As shown in Figure 3, CorrSteer performs effectively even with around 100 training samples, with no substantial improvements beyond 4,000 samples, making it practical for quick deployment.
+The high variance observed in CorrSteer-A for smaller datasets like GSM8K (1,000 samples) and HarmBench (108 samples) suggests that approximately 4,000 samples are recommended for stable performance.
+5.3 Side Effect Trade-offs
+To quantify trade-offs, we measure Side Effect Ratio (SER), capturing the proportion of harmful to beneficial side effects.
+Table 2 and Figure 4 show CorrSteer consistently lowers SER compared to fine-tuning while preserving accuracy and highlights CorrSteer-P as the most balanced strategy, while CorrSteer-S minimizes SER in safety-critical tasks such as HarmBench and BBQ.
+Table 2: Side Effect Ratio (SER) results for CorrSteer variants on Gemma-2 2B across eight benchmarks.
+SER quantifies unintended side effects introduced by steering interventions, with values closer to 0 indicating stronger performance/safety preservation.
+CorrSteer-S CorrSteer-P CorrSteer-A Fine-tuning
+Task SER NEG POS SER NEG POS SER NEG POS SER NEG POS
+MMLU 0.36 11 20 0.17 249 286 0.20 264 299 0.41 1108 1616
+MMLU-Pro 0.42 8 11 0.42 30 41 0.42 39 54 0.46 357 418
+GSM8K 0.56 20 16 0.67 31 15 0.52 63 59 0.65 213 116
+BBQ Ambig 0.00 0 658 0.00 0 1532 0.08 53 649 - - -
+BBQ Disambig 0.17 45 59 0.15 111 164 0.26 65 112 - - -
+HarmBench 0.25 2 6 0.14 4 24 0.04 3 67 - - -
+SimpleQA 0.17 1 5 0.19 3 13 0.35 6 11 - - -
+XSTest 0.33 7 10 0.52 7 10 0.47 14 5 - - -
+Figure 4: SER comparison between different CorrSteer variants for Gemma-2 2B.
+5.4 Effect of Pooling Strategies on Steering
+As discussed in Section 4, pooling determines how SAE activations are aggregated across tokens.
+To empirically validate these design choices, we conducted controlled experiments comparing mean-pooling, all-token pooling, and max-pooling across representative benchmarks covering reasoning (GSM8K), bias (BBQ), and safety (HarmBench, XSTest).
+The comparison is summarized in Table 5.
+Our results reveal clear trends.
+On multi-token generation tasks, mean-pooling severely degrades performance (e.g., HarmBench: 0.00%, XSTest: 53.65%), confirming that averaging dilutes the sparse but informative signals needed for steering.
+All-token pooling similarly underperforms, suggesting that aggregating contributions from every token introduces substantial noise.
+By contrast, max-pooling consistently outperforms alternatives across tasks, capturing salient activations while filtering out irrelevant ones.
+These findings validate our choice of max-pooling as the default aggregation strategy for correlation-based feature selection and steering.
+We note one exception: for long reasoning tasks such as GSM8K, mean-pooling provides more stable coefficient scaling, as max-pooling can produce excessively large coefficients when applied to every generated token.
+A more detailed ablation study, including per-task breakdowns and coefficient-scaling effects, is provided in Appendix A.4.
+5.5 Feature Interpretability and Transferability
+Selected features align with task requirements: structured output features dominate multiple-choice benchmarks (MMLU, BBQ), refusal-related features drive safety improvements (HarmBench), and domain-specific semantics contribute to specialized evaluations.
+Post-hoc analysis via Neuronpedia descriptions further supports their semantic relevance.
+Feature activation frequencies vary across tasks, with performance gains tracking activation dynamics (Appendix 7).
+Mathematical features also emerge across tasks, including bias and safety, consistent with findings that math-oriented pre-training improves broad accuracy (Shao et al., 2024).
+For BBQ features in LLaMA-3.1 8B (full list in Appendix A.8.2), positively correlated features emphasize neutrality and balance:
+  • L15/25166 themes of neutrality and balance in discourse (coeff: 0.259, corr: 0.433)
+  • L25/10753 expressions of perception or belief in social dynamics (coeff: 1.147, corr: 0.428)
+These results suggest that task-specific semantic features contribute more to accuracy than general recognition features.
+Our ablation further confirms that SAE-based sparse feature selection outperforms raw activation steering (Table 7).
+Feature Set Transferability.
+Cross-task experiments show that MMLU features transfer well, outperforming task-specific features on BBQ Ambig and performing comparably on MMLU-Pro (Table 8).
+This suggests that certain feature sets capture reasoning patterns shared by multiple-choice benchmarks.
+Task-Level Circuit and Spurious Correlation.
+CorrSteer’s multi-layer steering relates to circuit discovery research (Olah et al., 2020; Elhage et al., 2021).
+While prior work isolates task-specific circuits (Conmy et al., 2023; Marks et al., 2025; Ameisen et al., 2025; Lindsey et al., 2025; Sun, 2025), our steering vectors act as additive subgraphs across layers.
+Restricting feature selection to activations during answer generation reduces spurious correlations, and interventions consistently improve performance (Table 1, Table 3), indicating the effectiveness of the selected feature sets.
+Correlation for Selection, Intervention for Causality.
+CorrSteer employs correlation as a feature selection mechanism, then establishes causal relationships through direct steering interventions within the controlled LLM computational graph.
+Unlike spurious correlations with uncontrolled confounding variables, correlations within LLM circuits can be directly validated through residual stream intervention.
+The consistent performance improvements across tasks (Table 1, Table 3) demonstrate causal influence of selected features.
+6 Conclusion and Limitations
+This work introduces CorrSteer, a fully automated correlation-driven pipeline that enables generation-time discovery of steering-effective SAE features.
+By correlating task performance with specific activation patterns during inference, our method eliminates the dependency on contrastive datasets that has limited prior steering approaches.
+Across eight benchmarks, CorrSteer achieves substantial improvements in question answering, bias mitigation, and safety with minimal computational overhead and reduced side effects, while revealing semantically aligned steering circuits across multiple layers.
+By leveraging SAE’s inherently linear architecture where features are designed to be linearly combined, this design yields interpretable feature combinations without parameter modification, establishing linear correlation as a principled foundation for mechanistic interpretability.
+Despite these advances, limitations remain.
+The fundamental constraint of steering vectors lies in their static nature, which prevents adaptation to dynamic model behaviors.
+This particularly affects tasks requiring contextual adaptation or multi-step reasoning, where static steering cannot adequately handle the conditional nature of problem-solving processes.
+Furthermore, our correlation-based approach exhibits increased performance variance with smaller sample sizes, and the task-optimized features show limited cross-task transferability beyond single-token generation scenarios.
+Appendix A Appendix
+A.1 Implementation Details
+Feature Extraction:
+Feature selection employs 4,000 samples across all datasets.
+For fair comparison, the same samples are used for training fine-tuning models.
+When datasets contain fewer than 4,000 samples, we use all available data.
+For datasets without predefined train/validation/test splits, we allocate 27% for training, 3% for validation, and 70% for testing.
+GSM8K uses 1,000 samples for feature selection with 50 samples reserved for validation.
+Feature Steering:
+Steering interventions are applied at the pre-execution stage of each transformer layer.
+The first layer is excluded from steering as the token embedding layer predominantly contains spurious correlations unrelated to the target tasks.
+Evaluation Metrics: For multiple-choice tasks (MMLU, MMLU-Pro, BBQ), exact match accuracy is used under zero-shot evaluation.
+All results are reported as mean ± standard deviation across multiple random seeds for statistical robustness: 5 seeds for most tasks, 3 seeds for GSM8K.
+For Gemma-2 2B, the non-steered MMLU performance (52.23%) is lower than the Gemma-2 2B-IT 5-shot result (56.1%) reported in the original Gemma paper due to the zero-shot setting and lack of in-context learning examples.
 For safety benchmarks, 1 - ASR (Attack Success Rate) is computed using a small refusal-detection language model.
@@
 Detailed configurations are provided in Appendix A.1.
-Pooling Strategy: Two pooling strategies are available for coefficient and correlation calculations: max-pooling and mean-pooling.
-For multi-token generation tasks, max-pooling consistently outperforms mean-pooling, as empirically demonstrated in Table 5, likely due to its better capture of peak feature activations relevant to task success.
-However, for coefficient calculation in longer generation tasks such as GSM8K reasoning, mean-pooling is preferred as max-pooling produces excessively large coefficient values.
-Applying these large coefficients to every generated token degrades performance, leading to the adoption of mean-pooling for reasoning tasks.
-4.2 SER Comparison
-This study aims to demonstrate that CorrSteer can improve benchmark performance while maintaining low side effects.
-Using the defined SER, CorrSteer’s SER is compared against fine-tuning for a question-answering dataset.
-Additionally, the SER is compared when allowing or enforcing negatively correlated features, supporting the claim of SAE’s positive-tinted behavior.
-Finally, the SER is compared when pooling on every token rather than the inference-time generated token.
-4.3 Feature Interpretability Analysis
-The analysis is enhanced by incorporating feature descriptions from Neuronpedia, providing semantic interpretations of selected features.
-Safe/unsafe tendency analysis is conducted, along with task-wise and feature-level interpretations.
-This interpretability analysis validates that our correlation-based selection identifies semantically meaningful and task-relevant features, supporting the causal hypothesis underlying our approach.
-Table 1: Performance comparison between baseline and CorrSteer variants across BBQ, MMLU, MMLU-Pro, GSM8K, SimpleQA, and XSTest on Gemma 2 2B. Results show accuracy (%) for all tasks.
-Task Baseline CorrSteer-1 CorrSteer-P CorrSteer-A Fine-tuning
-BBQ Ambig 59.10 62.38 66.65 62.08 -
-BBQ Disambig 75.42 75.70 77.04 76.53 -
-HarmBench 44.64 46.07 51.79 67.50 -
-MMLU 52.23 52.82 55.83 56.32 55.85
-MMLU-Pro 30.30 30.44 30.82 31.01 33.16
-GSM8K 54.74 54.44 53.53 54.44 47.38
-SimpleQA 3.63 3.76 3.96 3.80 -
-XSTest 86.35 86.67 86.03 87.30 -
-Table 2: Performance comparison between baseline and CorrSteer variants across BBQ, MMLU, MMLU-Pro, HarmBench, SimpleQA, and XSTest on LLaMA 3.1 8B. Results show accuracy (%) for all tasks.
-Task Baseline CorrSteer-1 CorrSteer-P CorrSteer-A
+Fine-tuning
+Fine-tuning hyperparameters are determined through empirical experimentation across tasks and dataset sizes.
+Fine-tuning is performed using AdamW optimizer with learning rate 1e-5 (reduced to 5e-6 for small datasets <2000 samples), weight decay 0.01, and gradient clipping at norm 1.0.
+The training schedule includes 3% warmup steps followed by cosine annealing decay.
+Training proceeds for one epoch with 4,000 samples, using exact target supervision where prompt tokens are masked with -100 labels and only target spans contribute to the loss.
+A.2 Generation Benchmark Results
+Evaluation Models: Two specialized models are employed for evaluation.
+The DistillRoBERTa model [^1] https://huggingface.co/datasets/huggingface-community/distill-roberta-toxicity-r1 is used to identify the rejection of harmful requests, while the BERT STS model [^2] https://huggingface.co/datasets/HuggingFaceH4/stsb_multi_mt is used for matching generated answers against expected responses.
+A.3 Additional Results
+Figure 5: Benchmark performance of CorrSteer variants compared with the non-steered model on LLaMA-3.1 8B.
+Table 3: Performance comparison between non-steered model and CorrSteer variants across BBQ, MMLU, MMLU-Pro, HarmBench, SimpleQA, and XSTest on LLaMA-3.1 8B.
+Results show accuracy (%) under zero-shot evaluation (single-shot for BBQ).
+Task Non-steered Corrsteer-S CorrSteer-P CorrSteer-A
 BBQ Ambig 83.97 83.98 87.10 86.83
@@
 XSTest 61.27 62.22 62.22 58.41
-5 Results
-5.1 Generation Benchmark Performance
-Table 1 and Table 2 present comprehensive results across our evaluation benchmarks.
-CorrSteer demonstrates consistent improvements across most tasks, including question answering, bias mitigation, and safety benchmarks, with the exception of the GSM8K reasoning task.
-This pattern suggests that our method enhances model adherence to static task requirements without introducing dynamic behaviors.
-Method Comparison
-In most cases, CorrSteer-A and CorrSteer-P demonstrate the highest performance, with CorrSteer-P showing particular dominance in LLaMA 3.1 8B.
-This observation is attributed to the less disentangled nature of Llama Scope features from superposition, which necessitates more aggressive pruning.
-Safety Benchmarks:
-For HarmBench, selected features enhance the model’s ability to refuse harmful requests, achieving a 22.9% improvement.
-However, for XSTest, improvement is limited to 1% due to the overrefusal characteristics of the benchmark.
-This is an expected result due to the static nature of CorrSteer, which hinders the ability to clearly distinguish between benign and harmful requests.
-Table 3: Safety Evaluation Rate (SER) analysis for CorrSteer variants across different benchmarks on Gemma 2 2B. SER values closer to 0 indicate better safety performance.
-CorrSteer-1 CorrSteer-P CorrSteer-A Fine-tuning
-Task SER neg pos SER neg pos SER neg pos SER neg pos
-BBQ Ambig 0.000 0 658 0.000 0 1532 0.076 53 649 - - -
-BBQ Disambig 0.167 45 59 0.153 111 164 0.257 65 112 - - -
-HarmBench 0.250 2 6 0.143 4 24 0.043 3 67 - - -
-MMLU 0.355 11 20 0.172 249 286 0.202 264 299 0.407 1108 1616
-MMLU-Pro 0.421 8 11 0.423 30 41 0.419 39 54 0.461 357 418
-GSM8K 0.556 20 16 0.674 31 15 0.516 63 59 0.647 213 116
-SimpleQA 0.167 1 5 0.188 3 13 0.353 6 11 - - -
-XSTest 0.333 7 10 0.520 7 10 0.467 14 5 - - -
-Table 4: Safety Evaluation Rate (SER) analysis for CorrSteer variants on Llama models across different benchmarks. SER values closer to 0 indicate better safety performance.
-CorrSteer-1 CorrSteer-P CorrSteer-A
+Task-Specific Analysis
+MMLU: The global method selects features related to structured output formatting, addressing Gemma-2 2B’s tendency to generate tokens outside the required A/B/C/D options.
+Post-steering, this hallucination issue is largely resolved.
+MMLU-Pro: A similar issue occurs more severely due to the 10 options in MMLU-Pro.
+Constrained decoding, which samples tokens exclusively from available options, is applied to improve the model’s authentic capability, resulting in performance that remains higher than the non-steered model, with CorrSteer-A achieving maximum performance.
+BBQ: Similar improvements in format adherence are observed, with selected features promoting appropriate response structure.
+Figure 6: SER comparison across datasets between different CorrSteer variants on LLaMA-3.1 8B.
+Table 4: Side Effect Ratio (SER) analysis for CorrSteer variants on LLaMA-3.1 8B across different benchmarks. SER values closer to 0 indicate better safety performance.
+Corrsteer-S CorrSteer-P CorrSteer-A
 Task SER neg pos SER neg pos SER neg pos
@@
 XSTest 0.412 7 10 0.412 7 10 0.737 14 5
-Observations:
-Results for both Gemma-2 2B and LLaMA 3.1 8B, presented in Table 1 and Table 2, demonstrate consistent patterns, with CorrSteer variants showing systematic improvements.
-Multi-layer Superiority: Simultaneous multi-layer steering approaches, such as CorrSteer-A and CorrSteer-P, consistently outperform the single-layer CorrSteer-1 approach across all benchmarks, consistent with findings from other researchers (Liu et al., 2024; Zhao et al., 2025).
-This observation suggests that enabling features from different layers to compete globally yields superior task-relevant selections compared to layer-wise optimization.
-Limited Factuality Impact: SimpleQA shows minimal improvement, confirming that CorrSteer enhances task adherence without introducing external factual knowledge.
-This is a desirable property, as it indicates the method improves model behavior rather than injecting information not present in the original model.
-Comparison with Fine-tuning: CorrSteer demonstrates competitive performance while maintaining significantly lower side effect rates compared to fine-tuning.
-On MMLU, CorrSteer-A achieves higher accuracy (56.32%) than fine-tuning (55.85%) with substantially lower SER (0.202 vs 0.407).
-Similarly, on GSM8K, CorrSteer variants outperform fine-tuning in accuracy while maintaining lower side effect rates across all tasks.
-5.2 Feature Analysis
-Selected features demonstrate varying degrees of alignment with task requirements: while some benchmarks show consistent alignment through structured output features for multiple-choice tasks, refusal-related features for safety benchmarks, and task-specific semantic features for domain-specific evaluations, others exhibit less consistent patterns.
-The interpretability of selected features, validated through Neuronpedia descriptions, provides confidence in the semantic relevance of our correlation-based selection process.
-Notably, feature activation frequencies vary significantly across tasks, with performance improvements correlating with distinct activation patterns (Appendix 8).
-Analysis of selected features reveals semantically meaningful patterns aligned with task requirements:
-Safety-related Features: HarmBench and BBQ demonstrate substantial improvements through selected safety-related features that enhance neutrality and refusal behavior.
-This interpretability analysis is covered in detail in Section 6.1.
-5.3 SER Analysis
-Table 3 presents the Safety Evaluation Rate analysis, demonstrating CorrSteer’s ability to minimize side effects while improving task performance.
-Figure 5: SER comparison between different CorrSteer variants for Gemma 2 2B.
-6 Discussion
-This work establishes a viable and efficient approach for SAE-based steering, providing effective control across diverse applications.
-The interpretable feature combinations that yield performance improvements support the hypothesis that linear correlation serves as a meaningful unit for interpretable AI capabilities.
-6.1 Feature Inspection
-g A notable finding is that features selected by CorrSteer demonstrate task-relevant patterns that align with theoretical expectations Math-related features are consistently discovered across all tasks, proving beneficial even for bias mitigation and safety tasks.
-This universal correlation between mathematical features and accuracy aligns with DeepSeekMath (Shao et al., 2024)’s findings, where further pre-training on math-focused corpora yielded performance improvements across diverse tasks.
-For BBQ features in LLaMA 3.1 8B, we observe:
-  • L15/25166 themes of neutrality and balance in discourse (coeff: 0.259, corr: 0.433)
-  • L25/10753 expressions of perception or belief in social dynamics (coeff: 1.147, corr: 0.428)
-While bias-related features were expected for the BBQ benchmark, these neutrality-focused features demonstrate high positive correlation.
-Conversely, explicit bias-related and choice-making features exhibit negative correlations:
-  • L8/8123 questions that ask for truthfulness or correctness regarding options or statements (coeff: 3.725, corr: -0.133)
-  • L17/9134 choice-related phrases and expressions of preference (coeff: 2.379, corr: -0.451)
-  • L19/15745 phrases related to decision-making and choice, particularly in the context of parenting and social interactions (coeff: 9.740, corr: -0.464)
-These findings suggest that task-specific induced features contribute more to sample accuracy than meta-cognitive recognition features.
-Our ablation study further demonstrates that SAE-based sparse feature selection consistently outperforms raw activation steering across all evaluated tasks (Table 7).
-6.2 Feature Set Transferability
-The transferability of CorrSteer feature sets is evaluated across MMLU, MMLU-Pro, and BBQ benchmarks.
-Interestingly, our cross-task experiments reveal that MMLU features outperform task-specific features on BBQ Ambig and achieve comparable performance on MMLU-Pro, suggesting that some feature sets capture more generalizable reasoning patterns (Table 8).
-This transferability is attributed to their shared multiple-choice format, which requires similar structural feature patterns.
-6.3 Task-Level Circuit
-CorrSteer’s multi-layer approach relates to neural network circuit discovery research (Olah et al., 2020; Elhage et al., 2021).
-While emerging works focus on discovering task-specific circuits (Conmy et al., 2023; Marks et al., 2025; Ameisen et al., 2025; Lindsey et al., 2025; Sun, 2025), our steering vectors that work simultaneously across layers can be conceptualized as additive subgraphs of optimized task circuits, though they lack explicit interpretation of interactions and causality.
-6.4 Side Effect Ratio
-The primary challenge in AI steering concerns robustness for industrial applications, necessitating precise control mechanisms.
-Direct steering at each layer without updating original parameters based on token prediction distributions represents a key approach to minimize side effects.
-Theoretically, separating steering vectors across different activation spaces minimizes mutual interference in superpositioned states (Elhage et al., 2022).
-6.5 Pooling Strategy Analysis
-The results reveal interesting patterns across different pooling strategies (Table 5).
-Mean-pooling shows severe degradation on multi-token generation tasks (HarmBench: 0.00%, XSTest: 53.65%) where responses require multiple tokens.
-All-token pooling shows degraded performance on every task compared to max-pooling’s inference-time aggregation.
-This suggests that max-pooling better captures the critical activations needed for effective steering across all task types, while all-token pooling may introduce noise by including irrelevant token positions, and mean-pooling dilutes important signals by averaging across all tokens in longer sequences.
-6.6 Computational Efficiency
-Our streaming correlation implementation achieves O(1) memory complexity with respect to training set size, making the approach scalable to large datasets.
-The pipeline exhibits computational efficiency with minimal sample requirements (200-400) as demonstrated in Figure 4 and completes feature extraction within minutes.
-Static feature sets and coefficients at inference time eliminate SAE dependency during deployment.
-6.7 Practical Applicability
-CorrSteer operates as an auxiliary mechanism that captures correlations difficult to detect through supervised fine-tuning, applicable with minimal side effects even after fine-tuning.
-The automated pipeline enables rapid deployment across tasks and domains without hyperparameter tuning.
-The method shows effectiveness across two model families, as demonstrated on both Gemma-2 2B and LLaMA-3.1 8B models.
-Importantly, our approach has broader implications for AI safety, as demonstrated by its effectiveness in both bias mitigation and amplification (Table 9), highlighting the need for responsible deployment of such steering capabilities.
-7 Conclusion
-This work introduces CorrSteer, a fully automated pipeline for SAE-based language model steering that leverages correlation analysis.
-Our approach addresses key limitations of existing SAE steering methods by identifying task-relevant features without requiring manual feature exploration or contrastive datasets.
-Experimental validation across diverse benchmarks demonstrates CorrSteer’s effectiveness, consistently improving performance on question answering, bias mitigation, and safety evaluation tasks.
-Selected features for safety, mathematical, and refusal-related tasks reveal the underlying objectives and required capabilities that drive task performance.
-Future Work
-Several promising directions emerge from this work: Prompt Engineering Comparison: Future studies should compare CorrSteer with prompt engineering approaches, as prompt-based methods are expected to exhibit higher SER due to their less targeted intervention mechanisms.
-Dynamic Steering for Reasoning: The performance degradation observed in GSM8K reasoning tasks suggests the need for dynamic steering approaches that can adapt to the sequential nature of mathematical problem-solving, moving beyond static feature interventions.
-Orthogonal Feature Projection: To further minimize side effects, future work could explore feature filtering techniques that project out components already activated in baseline features before applying steering, potentially reducing interference with existing model capabilities.
-Acknowledgments
-The authors thank the teams behind Gemma Scope and LLaMA Scope for providing high-quality SAE releases that enabled this research.
-The authors also acknowledge Neuronpedia for providing automated feature descriptions that enhanced our interpretability analysis.
-Limitations
-The fundamental limitation of steering vectors is their static nature, which prevents adaptation to dynamic model behaviors.
-This constraint particularly affects reasoning tasks like GSM8K, where static steering cannot adequately handle the sequential nature of mathematical problem-solving.
-Our evaluation focuses primarily on discriminative and short-form generation tasks; long-form generation and creative tasks may require different approaches or modifications to our method.
-Appendix A Appendix
-A.1 Implementation Details
-Feature Extraction:
-Feature selection employs 4,000 samples across all datasets.
-For fair comparison, the same samples are used for training fine-tuning baselines.
-When datasets contain fewer than 4,000 samples, we use all available data.
-For datasets without predefined train/validation/test splits, we allocate 27% for training, 3% for validation, and 70% for testing.
-GSM8K uses 1,000 samples for feature selection with 50 samples reserved for validation.
-Feature Steering:
-Steering interventions are applied at the pre-execution stage of each transformer layer.
-The first layer is excluded from steering as the token embedding layer predominantly contains spurious correlations unrelated to the target tasks.
-Fine-tuning
-Fine-tuning is performed using AdamW optimizer with learning rate 1e-5 (reduced to 5e-6 for small datasets <2000 samples), weight decay 0.01, and gradient clipping at norm 1.0.
-The training schedule includes 3% warmup steps followed by cosine annealing decay.
-Training proceeds for one epoch with 4,000 samples, using exact target supervision where prompt tokens are masked with -100 labels and only target spans contribute to the loss.
-A.2 Generation Benchmark Results
-Evaluation Models: Two specialized models are employed for evaluation.
-The DistillRoBERTa model [^1] https://huggingface.co/datasets/huggingface-community/distill-roberta-toxicity-r1 is used to identify the rejection of harmful requests, while the BERT STS model [^2] https://huggingface.co/datasets/HuggingFaceH4/stsb_multi_mt is used for matching generated answers against expected responses.
-A.3 Additional Results
-Figure 6: Benchmark performance of CorrSteer variants compared with the default baseline on LLaMA 3.1 8B.
-Task-Specific Analysis
-MMLU: The global method selects features related to structured output formatting, addressing Gemma-2’s tendency to generate tokens outside the required A/B/C/D options.
-Post-steering, this hallucination issue is largely resolved.
-MMLU-Pro: A similar issue occurs more severely due to the 10 options in MMLU-Pro.
-Constrained decoding, which samples tokens exclusively from available options, is applied to improve the model’s authentic capability, resulting in performance that remains higher than baseline, with CorrSteer-A achieving maximum performance.
-BBQ: Similar improvements in format adherence are observed, with selected features promoting appropriate response structure.
-Figure 7: SER comparison across datasets between different CorrSteer variants on LLaMA 3.1 8B.
 Feature Frequency Analysis
@@
 Even for tasks with lower feature frequencies, CorrSteer maintains its advantage by preserving low SER values.
-Figure 8: Frequency of activation samples across layers of Gemma 2 2B for SimpleQA (left) and HarmBench (right) tasks.
+Figure 7: Frequency of activation samples across layers of Gemma-2 2B for SimpleQA (left) and HarmBench (right) tasks.
 A.4 Ablation Study
@@
 We also evaluate alternative pooling strategies including mean-pooling and all-token pooling for feature activation aggregation. The comparison results are presented in Table 5.
-Table 5: Pooling strategy comparison on Gemma 2 2B using CorrSteer-A.
+Table 5: Pooling strategy comparison on Gemma-2 2B using CorrSteer-A.
 For single-token generation such as MMLU, MMLU-Pro, and BBQ, mean-pooling naturally achieves identical performance to max-pooling since only one token is generated, while all-token pooling shows degraded performance.
 Mean-pooling shows severe degradation on multi-token generation tasks, demonstrating the superiority of max-pooling.
-Task Baseline Max-pooling Mean-pooling All-token
+Task Non-steered Max-pooling Mean-pooling All-token
 MMLU 52.23 56.32 56.32 52.91
@@
 To validate our design choice of using only positively correlated features, we conduct ablation experiments using negatively correlated features for steering.
-We compare two approaches: single-layer negative steering (CorrSteer-1 with negative features) and multi-layer negative steering (CorrSteer-A with negative features).
-Table 6: Performance comparison between positive and negative correlation feature steering on Gemma 2 2B.
+We compare two approaches: single-layer negative steering (Corrsteer-S with negative features) and multi-layer negative steering (CorrSteer-A with negative features).
+For Gemma-2 2B, negatively correlated features are bias- and choice-related:
+  • L8/8123 questions asking for correctness of options (coeff: 3.725, corr: -0.133)
+  • L17/9134 choice-related phrases and preferences (coeff: 2.379, corr: -0.451)
+  • L19/15745 decision-making and choice expressions in social contexts (coeff: 9.740, corr: -0.464)
+Table 6: Performance comparison between positive and negative correlation feature steering on Gemma-2 2B.
 Negative correlation features consistently show poor performance, validating our positive-only approach.
-Task Baseline CorrSteer-A Negative-1 Negative-G
+Task Non-steered CorrSteer-A Negative-S Negative-G
 MMLU 52.23 56.32 52.24 49.45
@@
 The results demonstrate that negative correlation features provide minimal improvement in single-layer steering and often cause severe performance degradation in multi-layer steering.
-Notably, MMLU-Pro drops to 0.66% and BBQ Disambig to 12.15% with negative multi-layer steering, confirming that negative correlations often represent spurious patterns rather than causal relationships.
+Notably, MMLU-Pro drops to 0.66% and BBQ Disambig to 12.15% with negative multi-layer steering, confirming that negative correlations often represent spurious patterns rather than causal relationships, as negative activations are often unrelated noise (Joseph Bloom, 2024).
 Additionally, combining positive and negative features simultaneously yields inferior performance compared to positive-only selection.
 This validates our approach of using only positively correlated features, which aligns with the non-negative nature of SAE activations.
-Table 7: Performance comparison between raw activation steering and SAE-decoded steering on Gemma 2 2B.
+Table 7: Performance comparison between raw activation steering and SAE-decoded steering on Gemma-2 2B.
 Decoding adds SAE decoder bias term for the first layer, while Decoding-A adds multi-layer feature directions as CorrSteer-A.
-Task Baseline Raw Activation Decoding-1 Decoding-A CorrSteer-A
+Task Non-steered Raw Activation Decoding-S Decoding-A CorrSteer-A
 MMLU 52.23 49.85 55.38 54.38 56.32
@@
 The results demonstrate a clear performance hierarchy: CorrSteer-A > SAE Decoding > Raw Activation across all evaluated tasks, which is explainable by Superposition Hypothesis (Elhage et al., 2022).
-One exception occurred in BBQ Disambig, where Decoding-1 shows better performance than CorrSteer-A.
-However, Decoding-1 failed to show robustness across benchmarks, frequently degrading performance while CorrSteer-A shows consistent performance across all tasks.
+One exception occurred in BBQ Disambig, where Decoding-S shows better performance than CorrSteer-A.
+However, Decoding-S failed to show robustness across benchmarks, frequently degrading performance while CorrSteer-A shows consistent performance across all tasks.
 SAE Decoder Bias
@@
 This analysis provides insights into the generalizability of task-specific feature sets.
-Table 8: Cross-task feature transferability results on Gemma 2 2B.
+Table 8: Cross-task feature transferability results on Gemma-2 2B.
 Features selected from source tasks (rows) are applied to target tasks (columns).
-Results show accuracy (%) with baseline performance in parentheses.
-MMLU-Pro results do not use constrained decoding, achieving 17.56% compared to unconstrained baseline (14.00%).
+Results show accuracy (%) with non-steered model performance in parentheses.
+MMLU-Pro results do not use constrained decoding, achieving 17.56% compared to unconstrained non-steered model (14.00%).
 Source → Target MMLU MMLU-Pro BBQ Disambig BBQ Ambig
@@
 Table 9: Bias steering effectiveness across different demographic categories on EMGSD dataset. Mitigation reduces bias scores, while amplification increases them.
-Mitigation Amplification
-Category Baseline CorrSteer Biased CorrSteer
+Mitigation (Fairness $\uparrow$) Amplification (Bias $\uparrow$)
+Category Non-steered CorrSteer Biased CorrSteer
 Gender 0.177 0.616 0.897 0.922
@@
 Results demonstrate that correlation-selected features provide effective steering control across all demographic categories (Table 9).
-Our steering approach effectively reduces bias, with steered outputs showing substantially lower bias scores compared to biased baselines, supporting the generalizability of our approach across different domains and SAE training paradigms.
+For mitigation, CorrSteer surpasses the non-steered model across categories by improving fairness scores.
+For amplification, CorrSteer generally increases bias relative to the biased non-steered model, with the LGBTQ+ row as an exception to be audited.
 A demonstration of our bias mitigation results is available at https://huggingface.co/spaces/seonglae/CorrSteer, showcasing real-time steering capabilities.
-A.7 Complete Feature Lists
+A.7 Framework Implications
+CorrSteer leverages generation-time activations for multi-token, multi-layer SAE-based steering, and our experiments are enabled by Gemma Scope (Lieberum et al., 2024) and LLaMA Scope (He et al., 2024), the only open releases providing SAEs across all residual stream layers.
+The proposed framework demonstrates the practical utility of SAE in real-world LLM inference, addressing critical concerns such as safe reasoning, bias mitigation, and resistance to jailbreaking.
+This research demonstrates that SAE-based control mechanisms offer a promising direction for both understanding and improving LLM behavior.
+The framework’s ability to operate through an interpretable interface while maintaining or improving model performance suggests a concrete path toward safer, more transparent AI.
+A.8 Complete Feature Lists
 This section presents the complete feature lists for each task, showing the top-1 features aggregated from all layers.
@@
 Following each layer’s highest correlated feature, we include additional relevant features listed below.
-A.7.1 Gemma-2B
+A.8.1 Gemma-2B
 BBQ (Ambiguous)
-Figure: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of Gemma 2 2B.
+Figure 8: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of Gemma-2 2B.
   • L1/6088 specific formatting or structural elements within text, such as timestamps and code (coeff: 2.280, corr: 0.134)
@@
   • L19/15745 phrases related to decision-making and choice, particularly in the context of parenting and social interactions (coeff: 9.740, corr: -0.464)
-Figure: Top correlated features with BBQ ambig on frequency in each layer of Gemma 2 2B.
+Figure 9: Top correlated features with BBQ ambig on frequency in each layer of Gemma-2 2B.
 BBQ (Disambiguous)
-Figure: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of Gemma 2 2B.
+Figure 10: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of Gemma-2 2B.
   • L1/7001 code structure and elements in programming, particularly related to class and variable definitions (coeff: 2.126, corr: 0.114)
@@
   • L23/12048 terms related to racism and social injustice (coeff: 2.661, corr: 0.147)
-Figure: Top correlated features with BBQ disambig on frequency in each layer of Gemma 2 2B.
+Figure 11: Top correlated features with BBQ disambig on frequency in each layer of Gemma-2 2B.
 HarmBench
-Figure: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of Gemma 2 2B.
+Figure 12: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of Gemma-2 2B.
   • L1/9572 occurrences of the semicolon character (coeff: 5.206, corr: 0.692)
@@
   • L25/3912 negative sentiments or refusals (coeff: 57.431, corr: 0.711)
-Figure: Top correlated features with HarmBench on frequency in each layer of Gemma 2 2B.
+Figure 13: Top correlated features with HarmBench on frequency in each layer of Gemma-2 2B.
 MMLU
-Figure: Top correlated features with selected features from CorrSteer-P with MMLU on coefficient in each layer of Gemma 2 2B.
+Figure 14: Top correlated features with selected features from CorrSteer-P with MMLU on coefficient in each layer of Gemma-2 2B.
   • L1/13714 colons and semicolons used in lists or programming syntax (coeff: 0.403, corr: 0.140)
@@
   • L25/10388 phrases related to health-related actions and topics (coeff: 33.899, corr: 0.336)
-Figure: Top correlated features with MMLU on frequency in each layer of Gemma 2 2B.
+Figure 15: Top correlated features with MMLU on frequency in each layer of Gemma-2 2B.
 MMLU-Pro
-Figure: Top correlated features with selected features from CorrSteer-P with MMLU-Pro on coefficient in each layer of Gemma 2 2B.
+Figure 16: Top correlated features with selected features from CorrSteer-P with MMLU-Pro on coefficient in each layer of Gemma-2 2B.
   • L1/9317 phrases related to changes in social and organizational dynamics (coeff: 1.859, corr: 0.169)
@@
   • L25/8643 scientific terms and concepts related to biochemistry and cellular processes (coeff: 11.439, corr: 0.545)
-Figure: Top correlated features with MMLU-Pro on frequency in each layer of Gemma 2 2B.
+Figure 17: Top correlated features with MMLU-Pro on frequency in each layer of Gemma-2 2B.
 GSM8K
-Figure: Top correlated features with selected features from CorrSteer-P with GSM8K on coefficient in each layer of Gemma 2 2B.
+Figure 18: Top correlated features with selected features from CorrSteer-P with GSM8K on coefficient in each layer of Gemma-2 2B.
   • L1/13475 specific quantitative or statistical information (coeff: 9.936, corr: 0.251)
@@
   • L25/3469 technical aspects related to semiconductor devices and their manufacturing processes (coeff: 23.158, corr: 0.284)
-Figure: Top correlated features with GSM8K on frequency in each layer of Gemma 2 2B.
+Figure 19: Top correlated features with GSM8K on frequency in each layer of Gemma-2 2B.
 SimpleQA
-Figure: Top correlated features with selected features from CorrSteer-P with SimpleQA on coefficient in each layer of Gemma 2 2B.
+Figure 20: Top correlated features with selected features from CorrSteer-P with SimpleQA on coefficient in each layer of Gemma-2 2B.
   • L1/14904 references to Congress and legislative processes (coeff: 0.263, corr: 0.192)
@@
   • L3/12843 terms related to durability and long-lasting qualities (coeff: 0.219, corr: 0.178)
-  • L4/680 references to standards, particularly in legal and medical contexts (coeff: 0.055, corr: 0.194)
-  • L5/4460 references to legal cases and court rulings (coeff: 0.093, corr: 0.193)
-  • L6/9777 expressions of agreement or dissent and the context surrounding them (coeff: 0.153, corr: 0.192)
-  • L7/2431 phrases related to time management and constraints (coeff: 0.253, corr: 0.213)
-  • L8/14209 HTML coding elements and formatting commands (coeff: 0.097, corr: 0.194)
-  • L9/7856 terms related to penalties and scoring in sporting events (coeff: 0.257, corr: 0.196)
-  • L10/2446 terms related to health and legal matters (coeff: 0.350, corr: 0.177)
-  • L11/7954 legal terminology and references to court cases and proceedings (coeff: 0.164, corr: 0.194)
-  • L12/1495 phrases related to the duration and continuity of experiences over time (coeff: 0.208, corr: 0.195)
-  • L13/12119 references to various parameters and aspects within scientific or technical contexts (coeff: 0.148, corr: 0.194)
-  • L14/6355 references to India and its cultural context (coeff: 0.377, corr: 0.201)
-  • L15/4385 programming constructs related to class and method declarations in Java (coeff: 0.177, corr: 0.194)
-  • L16/7182 expressions of enthusiasm or amazement (coeff: 0.457, corr: 0.216)
-  • L17/6346 references to offices or organizational structures (coeff: 0.274, corr: 0.194)
-  • L18/5258 terms related to legal charges and prosecutions (coeff: 0.843, corr: 0.200)
-  • L19/4202 technical terms and concepts related to physical phenomena and their mathematical descriptions (coeff: 0.262, corr: 0.194)
-  • L20/6557 legal terms and phrases related to criminal charges and legal proceedings (coeff: 0.953, corr: 0.220)
-  • L21/13830 references to political leaders and government roles (coeff: 8.090, corr: 0.196)
-  • L22/15897 phrases related to international relations and cooperation, particularly in the context of political statements and actions (coeff: 0.648, corr: 0.237)
-  • L23/15190 terms related to health, well-being, and interventions for obesity and mental illness (coeff: 0.832, corr: 0.237)
-  • L24/15228 references to political figures and their actions (coeff: 2.043, corr: 0.222)
-  • L25/2531 expressions of political opinion regarding government spending and fiscal policies (coeff: 1.664, corr: 0.200)
-Figure: Top correlated features with SimpleQA on frequency in each layer of Gemma 2 2B.
-XSTest
-Figure: Top correlated features with selected features from CorrSteer-P with XSTest on coefficient in each layer of Gemma 2 2B.
-  • L1/4509 terms and concepts related to scientific and mathematical structures and functions (coeff: 1.940, corr: 0.333)
-  • L2/4679 financial metrics and forecasts related to stock performance (coeff: 1.584, corr: 0.301)
-  • L3/1326 legal terminology and references to statutes and claims (coeff: 4.088, corr: 0.256)
-  • L4/12152 references to geographical locations and their associated attributes (coeff: 0.961, corr: 0.259)
-  • L5/5939 terms related to signals and their coding in biological contexts (coeff: 3.391, corr: 0.248)
-  • L6/4376 numerical values and specific formatting related to data structures or coding (coeff: 6.713, corr: 0.259)
-  • L7/4886 representations of numerical data, particularly in scientific contexts (coeff: 3.074, corr: 0.286)
   • L8/10825 punctuation marks and special characters (coeff: 5.194, corr: 0.296)
@@
   • L25/7968 elements related to health assessments and metrics (coeff: 9.863, corr: 0.307)
-Figure: Top correlated features with XSTest on frequency in each layer of Gemma 2 2B.
-A.7.2 Llama-3.1-8B
+Figure 21: Top correlated features with XSTest on frequency in each layer of Gemma-2 2B.
+A.8.2 LLaMA-3.1-8B
 BBQ (Ambiguous)
-Figure: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of Llama 3.1 8B.
+Figure 22: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of LLaMA-3.1 8B.
   • L1/23207 phrases related to legal or regulatory frameworks (coeff: 0.463, corr: 0.111)
@@
   • L31/10728 auxiliary verbs and words indicating obligation or possibility (coeff: 1.530, corr: 0.239)
-Figure: Top correlated features with BBQ ambig on frequency in each layer of Llama 3.1 8B.
+Figure 23: Top correlated features with BBQ ambig on frequency in each layer of LLaMA-3.1 8B.
 BBQ (Disambiguous)
-Figure: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of Llama 3.1 8B.
+Figure 24: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of LLaMA-3.1 8B.
   • L1/5891 technical terms and references in programming and development contexts (coeff: 0.154, corr: 0.086)
@@
   • L31/18173 procedural commands and technical instructions related to software and settings (coeff: 1.440, corr: 0.188)
-Figure: Top correlated features with BBQ disambig on frequency in each layer of Llama 3.1 8B.
+Figure 25: Top correlated features with BBQ disambig on frequency in each layer of LLaMA-3.1 8B.
 HarmBench
-Figure: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of Llama 3.1 8B.
+Figure 26: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of LLaMA-3.1 8B.
   • L1/15747 repetitive phrases or expressions related to certainty or emphasis (coeff: 0.491, corr: 0.524)
@@
   • L31/31181 references to familial relationships and memorial details (coeff: 1.218, corr: 0.639)
-Figure: Top correlated features with HarmBench on frequency in each layer of Llama 3.1 8B.
+Figure 27: Top correlated features with HarmBench on frequency in each layer of LLaMA-3.1 8B.
 MMLU
-Figure: Top correlated features with selected features from CorrSteer-P with MMLU ambig on coefficient in each layer of Llama 3.1 8B.
+Figure 28: Top correlated features with selected features from CorrSteer-P with MMLU ambig on coefficient in each layer of LLaMA-3.1 8B.
   • L1/4557 specific numeric values and measurements related to instructions or guidelines (coeff: 0.695, corr: 0.094)
@@
   • L31/6722 instances of code-related syntax and formatting (coeff: 0.538, corr: 0.159)
-Figure: Top correlated features with MMLU on frequency in each layer of Llama 3.1 8B.
+Figure 29: Top correlated features with MMLU on frequency in each layer of LLaMA-3.1 8B.
 MMLU-Pro
-Figure: Top correlated features with selected features from CorrSteer-P with MMLU-Pro ambig on coefficient in each layer of Llama 3.1 8B.
+Figure 30: Top correlated features with selected features from CorrSteer-P with MMLU-Pro ambig on coefficient in each layer of LLaMA-3.1 8B.
   • L1/2403 specific numeric values and measurements related to instructions or guidelines (coeff: 0.286, corr: 0.216)
@@
   • L31/9926 numbers, particularly in relation to financial data and statistics (coeff: 10.348, corr: 0.280)
-Figure: Top correlated features with MMLU-Pro on frequency in each layer of Llama 3.1 8B.
+Figure 31: Top correlated features with MMLU-Pro on frequency in each layer of LLaMA-3.1 8B.
 SimpleQA
-Figure: Top correlated features with SimpleQA on frequency in each layer of Llama 3.1 8B.
+Figure 32: Top correlated features with SimpleQA on frequency in each layer of LLaMA-3.1 8B.
   • L1/28160 references to height, specifically focusing on the term "tall" (coeff: 1.580, corr: 0.454)
@@
   • L31/6171 references to artificial intelligence and its related concepts (coeff: 1.429, corr: 0.377)
-Figure: Top correlated features with SimpleQA on frequency in each layer of Llama 3.1 8B.
+Figure 33: Top correlated features with SimpleQA on frequency in each layer of LLaMA-3.1 8B.
 XSTest
-Figure: Top correlated features with XSTest on frequency in each layer of Llama 3.1 8B.
+Figure 34: Top correlated features with XSTest on frequency in each layer of LLaMA-3.1 8B.
   • L1/6754 references to studies and publications (coeff: 0.256, corr: 0.367)
@@
   • L31/2111 descriptions and features of software products (coeff: 1.614, corr: 0.276)
-Figure: Top correlated features with XSTest on frequency in each layer of Llama 3.1 8B.
+Figure 35: Top correlated features with XSTest on frequency in each layer of LLaMA-3.1 8B.